### v5

# Estimating Heterogeneous Causal Effects of High-Dimensional Treatments:

Abstract
Estimation of heterogeneous treatment effects is an active area of research. Most of the existing methods, however, focus on estimating the conditional average treatment effects of a single, binary treatment given a set of pre-treatment covariates. In this paper, we propose a method to estimate the heterogeneous causal effects of high-dimensional treatments, which poses unique challenges in terms of estimation and interpretation. The proposed approach finds maximally heterogeneous groups and uses a Bayesian mixture of regularized logistic regressions to identify groups of units who exhibit similar patterns of treatment effects. By directly modeling group membership with covariates, the proposed methodology allows one to explore the unit characteristics that are associated with different patterns of treatment effects. Our motivating application is conjoint analysis, which is a popular type of survey experiment in social science and marketing research and is based on a high-dimensional factorial design. We apply the proposed methodology to the conjoint data, where survey respondents are asked to select one of two immigrant profiles with randomly selected attributes. We find that a group of respondents with a relatively high degree of prejudice appears to discriminate against immigrants from non-European countries like Iraq. An open-source software package is available for implementing the proposed methodology.
Key words: causal inference, factorial design, mixture model, randomized experiment, regularized regression
1 Introduction
Over the past decade, a number of researchers have exploited modern machine learning algorithms and proposed new methods to estimate heterogeneous treatment effects using experimental data. They include tree-based methods (e.g., Imai and Strauss, 2011; Athey and Imbens, 2016; Wager and Athey, 2018; Hahn, Murray and Carvalho, 2020), regularized regressions (e.g., Imai and Ratkovic, 2013; Tian et al., 2014; Künzel et al., 2019), ensemble methods (e.g., van der Laan and Rose, 2011; Grimmer, Messing and Westwood, 2017), and frameworks that allow for the use of generic machine learning methods (e.g., Chernozhukov et al., 2019; Imai and Li, 2025). This methodological development, however, has largely been confined to settings with a single, binary treatment variable; some exceptions include a time-varying treatment (e.g., Almirall et al., 2014), and a relatively small number of treatments (e.g., Imai and Ratkovic, 2013).
In this paper, we estimate the heterogeneous effects of a high-dimensional treatment by analyzing the data from conjoint experiments, in which the number of possible treatment combinations exceeds the sample size. While the high dimensionality in treatment effect heterogeneity problems typically comes from the number of covariates or moderators, conjoint experiments provide an additional difficulty due to high dimensionality of treatment. We address the methodological challenge of effectively summarizing the complex patterns of heterogeneous treatment effects that are induced by the interactions among the treatments themselves as well as the interactions between the treatments and unit characteristics.
Methodological contributions.
We consider a common setting where researchers wish to use a small number of groups to summarize heterogeneous treatment effects and characterize these groups using several pre-treatment covariates (e.g., Chernozhukov et al., 2019; Imai and Li, 2025). We show that once researchers select the number of groups to be used for summarizing heterogeneous treatment effects, finding the maximally heterogeneous groups in terms of potential outcomes is equivalent to maximizing the likelihood function based on the latent group membership. Furthermore, modeling the conditional probability of an individual’s latent group membership using the moderators of interest yields maximally heterogeneous groups that are predicted well by these moderators.
A primary methodological challenge with high-dimensional treatments is characterizing both the interactions among a large number of treatment variables and their relationships with moderating covariates. Our methodology addresses this by finding maximally heterogeneous groups while characterizing the relationship between group membership and unit characteristics. Thus, it is possible to understand the types of units that are likely to exhibit similar treatment effect patterns.
Since optimizing over the latent group membership is difficult, we marginalize it out, leading to a mixture of experts model (e.g., Gormley and Frühwirth-Schnatter, 2019; Gupta and Chintagunta, 1994). We also develop estimation strategies by bringing together two previously disconnected literatures, one on mixture models and the other on sparsity-inducing penalties to fuse factor levels.
Empirical application.
Conjoint analysis is a popular survey experimental methodology in social sciences and marketing research (e.g., Hainmueller, Hopkins and Yamamoto, 2014; Rao, 2014). Conjoint analysis is a variant of factorial designs (Dasgupta, Pillai and Rubin, 2015) with a large number of factorial treatments—so large that typically not all possible treatments are observed. Under the most commonly used “forced-choice” design, respondents are asked to evaluate a pair of profiles whose attributes are randomly selected based on factorial variables with several levels.
In the specific experiment we reanalyze, the original authors used a conjoint analysis to measure immigration preferences by presenting each survey respondent with several pairs of immigrant profiles with varying attributes including education, country of origin, and job experience (Hainmueller and Hopkins, 2015). For each pair, the respondent was asked to choose one profile they prefer. The authors then analyzed the resulting response patterns to understand which immigrant characteristics play a critical role in forming the immigration preferences of American citizens.
In the methodological literature on factorial designs and conjoint analysis, researchers have focused on average marginal effects, which represent the average effect of one factor level relative to another level of the same factor averaging over the randomization distribution of the remaining factors (Hainmueller, Hopkins and Yamamoto, 2014; Dasgupta, Pillai and Rubin, 2015). Many empirical researchers use subgroup analysis to explore how these marginal effects depend on a small number of moderating covariates (e.g., Hainmueller and Hopkins, 2015; Newman and Malhotra, 2019).
Unfortunately, such an approach often results in low statistical power and may suffer from multiple testing problems (Liu and Shiraito, 2023). More fundamentally, by marginalizing other treatments, researchers may miss important interactions among treatments. Although some have explored the estimation of interaction effects (e.g., Dasgupta, Pillai and Rubin, 2015; Egami and Imai, 2019; De la Cuesta, Egami and Imai, 2022), few have investigated how to estimate heterogeneous treatment effects of high-dimensional treatments.
Moreover, there is even less prior research that models how the effects of high-dimensional treatments vary as a function of moderators. One exception is Robinson and Duch (2024) which uses a BART-based approach for conjoint experiments, but their heterogeneous effects of interest are different from ours (see Section 5.4 for comparison).
Related models.
To overcome this challenge, we develop a mixture of regularized logistic regression model under our general methodological framework of treatment effect heterogeneity with high-dimensional treatments. We combine and extend two distinct strands of methodological research. First, a growing literature explores regularization with high-dimensional factors, and their interactions, by fusing or grouping levels of factors together (e.g., Bondell and Reich, 2009; Post and Bondell, 2013; Stokell, Shah and Tibshirani, 2021). This methodology is well-suited to factorial experiments because it provides a natural way of interpreting empirical findings by identifying a set of factor levels that characterize distinct treatment effects (e.g., Egami and Imai 2019).
However, since our goal is to identify groups of individuals with heterogeneous effects, we use a mixture model that finds the maximally heterogeneous groups (see Section 3.2). Although the marketing literature has long applied mixture models to analyzing heterogeneity in conjoint experiments (e.g., Gupta and Chintagunta, 1994; Andrews, Ainslie and Currim, 2002), they focused on settings with low-dimensional treatments. In the high-dimensional setting, some combine mixture models with sparsity constraints (e.g., Khalili and Chen, 2007; Städler, Bühlmann and Van De Geer, 2010; Khalili, 2010), but these constraints are not designed to induce the fusion of factor levels that is essential in conjoint analysis.
Our model, therefore, synthesizes both of these approaches by using a finite mixture model with a prior that encourages fusing levels, while respecting the hierarchical structure—fusing main effects of factors only if their interactions are also fused (Yan and Bien, 2017). For efficient computation, we develop an Expectation-Maximization (EM) algorithm (Dempster, Laird and Rubin, 1977) by exploiting the representation of $\ell_{1}$ and $\ell_{2}$ penalties as a mixture of Gaussians (e.g., Figueiredo, 2003; Polson and Scott, 2011; Ratkovic and Tingley, 2017; Goplerud, 2021). We derive a tractable algorithm that adapts the latent overlapping group LASSO developed in sparse modeling to fusion required in factorial experiments.
The rest of the paper is organized as follows. In Section 2, we discuss the motivating application, which is a conjoint analysis of American citizens’ preferences regarding immigrant features. We also briefly describe a methodological challenge to be addressed. In Section 3, we present our proposed methodology. In Section 4, we show our method performs well in a realistic numerical simulation. In Section 5, we apply this methodology and reanalyze the data from the motivating conjoint analysis. Section 6 concludes with a discussion. The R package FactorHet (Goplerud, Pashley and Imai, 2025) can be used to implement our methodology and Goplerud, Imai and Pashley (2025) provides replication code for our application and simulations.
2 Motivating Application: Conjoint Analysis of Immigration Preferences
Our motivating application is a conjoint analysis of American immigration preferences. In this section, we introduce the experimental design and discuss the results of previous analyses that motivate our methodology for estimating heterogeneous treatment effects.
2.1 The Experimental Design
In an influential study, Hainmueller and Hopkins (2015) use conjoint analysis to estimate the effect of immigrant attributes on preferences for admission to the United States (Data are available at the AJPS Dataverse https://doi.org/10.7910/DVN/25505). The authors conduct an online survey experiment using a sample of 1,407 American adults. Each survey respondent assessed five pairs of immigrant profiles with randomly selected attributes. For each pair, a respondent was asked to choose which of the two immigrant profiles they preferred to admit to the United States.
The attributes of immigrant profiles used in this factorial experiment, with number of levels provided in parentheses, are gender (2), education (7), employment plans (4), job experience (4), profession (11), language skills (4), country of origin (10), reasons for applying (3), and prior trips to the United States (5). For completeness, these factors and their levels are reproduced as Table A1 of the Supplementary Material (Goplerud, Imai and Pashley, in-press). In total, there exist over 1.4 million possible profiles, implying more than $2\times 10^{12}$ possible comparisons of two profiles that are possible in the experiment. It is clear that with 1,407 respondents, even though each respondent performs five comparisons, not all possible profiles can be included. Thus, exploring treatment effect heterogeneity requires a methodological development that goes beyond the models used previously in the causal inference literature for binary treatments.
The levels of each factor variable were independently randomized to yield one immigrant profile. Randomization was subject to some restrictions such that profession and education factors result in sensible pairings (e.g., ruling out doctors with less than two-years of college education) and immigrants whose reason for applying is persecution must come from Iraq, Sudan, Somalia, or China. The ordering of attributes was also randomized for each respondent. The experiment additionally collected data on the respondents, including demographic information, partisanship, attitudes towards immigration, and ethnocentrism. A rating for each immigrant profile was also recorded, but that metric is not the focus of our analysis.
2.2 Heterogeneous Treatment Effects
Hainmueller and Hopkins (2015) conducted their primary analysis based on linear regression model where the unit of analysis is an immigrant profile (rather than a pair) and the outcome variable is an indicator for whether a given profile was chosen. The predictors of the model include the indicator variable for each immigrant attribute. The model also includes the interactions between education and profession, as well as between country of origin and reasons for applying, to account for the restricted randomization scheme mentioned above. Finally, the standard errors are clustered by respondent.
Figure 1: Estimated average marginal component effects of country of origin where the baseline is Germany, with effect estimates as given in Hainmueller and Hopkins (2015).
As formalized in Hainmueller, Hopkins and Yamamoto (2014), the regression coefficient represents the average marginal component effect (AMCE) of each attribute averaging over all the other attributes including those of the other profile in a given pair. Fig 1 reproduces the estimated overall AMCEs of country of origin where the baseline category is Germany. There is little country effect with the exception of Iraq, which negatively affects the likelihood of being preferred by a respondent.
Moderator Split
Education Any college education or no college education
Ethnocentrism Median ethnocentrism measure
Political party Republican or Democrat
Percent of foreign born workers in respondent’s industry High or low
Household income More or less than \$50,000
Fiscal exposure to immigration High or low
ZIP code demographics $<$ 5% immigrants, $>$ 5% immigrants (primarily from Latin America), or $>$ 5% immigrants (primarily not from Latin America)
Race/ethnicity White or non-white
Hispanic ethnicity Hispanic or non-hispanic
Ideology Liberal or conservative
Immigration attitudes Supports or does not support reducing immigration
Gender Male or female
Age Young or old
Table 1: List of subset analyses performed in Hainmueller and Hopkins (2015), listed by moderator and how it was split to form sub-groups.
Beyond the AMCEs, these authors and others including Newman and Malhotra (2019) have explored the heterogeneous treatment effects among respondents by conducting many sub-group analyses based on a number of respondent characteristics including partisanship and level of education. Table 1 shows all of the sub-group analyses performed by Hainmueller and Hopkins (2015) and how the respondents were broken up into groups. We find that 13 sub-group analyses were performed (excluding those used for robustness checks), with results from the first three (education, ethnocentrism, and political party) presented in the main paper. Of those three analyses, the authors find some evidence of heterogeneous effects of country of origin between subsets that differ on ethnocentrism, but little evidence of heterogeneity beyond this. The other 10 analyses can be found in their appendix, and the authors conclude for that participants responded similarly, in general, across those sub-groups.
Our goal is to build a methodology that enables one to more systematically explore heterogeneous treatment effects in conjoint experiments. Sub-group analyses like those conducted in the original analysis can be problematic for several reasons. First, the analyst must conduct a separate analysis for each moderator of interest, leading to multiple testing problem. Second, typically the moderators are dichotomized (or broken up into a small number of groups), requiring the analyst to decide how to split the data. Third, they are not amenable to exploration of how multiple moderators might work together to change outcomes.
To address these issues, one could include the moderators as covariates within the regression. However, if the goal is to provide estimated heterogeneous effects with straightforward interpretations, regressions with possibly complex interactions are not ideal. To estimate heterogeneous effects, we need to not only interact a large number of treatments, but we will have to further interact all main and interaction effects of treatments with the moderators. It is unclear how to best reduce the dimensionality of both the moderator and treatment space in a classic regression set up. It is also challenging to interpret the interactions from these models to understand the characteristics of units that lead to different treatment effect patterns.
In sum, researchers must parsimoniously characterize how a large number of possible treatment combinations interact with several key moderators of interest. The goal is to obtain estimates of heterogeneous effects and understand how the covariate distributions of units with different treatment effects differ. We now turn to our methodology which is designed to address these challenges and result in interpretable estimates.
3 Modeling Heterogeneous Effects of High-dimensional Treatments
We now describe the proposed methodology. To simplify the exposition, we focus on a general factorial design. This design corresponds to conjoint analysis with a single task per person, where there is only one profile assessed rather than a comparison of profiles, and complete randomization of all combinations of factor levels. Extensions to independent factor randomization and realistic conjoint analyses are immediate and will be discussed and applied in Section 5.
3.1 Set Up
Suppose that we have a simple random sample of $N$ units. Consider a factorial design with $J$ factors where each factor $j\in\{1,\cdots,J\}$ has $L_{j}\geq 2$ levels. The treatment variable for unit $i$, denoted by $\bm{T}_{i}$, is a $J$-dimensional vector of random variables, each of which represents the assigned level of the corresponding factor variable. For example, the $j$th element of this random vector $T_{ij}\in\{0,1,2,\ldots,L_{j}-1\}$ represents the level of factor $j$ which is assigned to unit $i$.
Following Dasgupta, Pillai and Rubin (2015), we define the potential outcome for unit $i$ as $Y_{i}(\bm{t})$ where $\bm{t}\in\mathcal{T}$ represents the realized treatment with $\mathcal{T}$ representing the support of the randomization distribution for $\bm{T}_{i}$. Then, the observed outcome is given by $Y_{i}=Y_{i}(\bm{T}_{i})$. The notation implicitly assumes no interference between units (Rubin, 1980). In this paper, for the sake of concreteness, we focus on the binary outcome $Y_{i}\in\{0,1\}$. Extensions to non-binary outcomes are straightforward. Lastly, we observe a vector of $p_{x}$ pre-treatment covariates for each unit $i$ and denote it by $\bm{X}_{i}$. All together, we observe $(Y_{i},\bm{T}_{i},\bm{X}_{i})$ for each unit $i$.
To illustrate the notation, consider a simplified version of our motivating example where each respondent $i$ observes a single immigrant profile and must decide whether to support that immigrant’s admission or not. Then, $\bm{T}_{i}$ is a vector indicating the level respondent $i$ sees for each of the nine immigrant attributes. The outcome variable $Y_{i}$ is an indicator for whether respondent $i$ chooses to support admission for the immigrant they are presented with. Lastly, $\bm{X}_{i}$ denotes a vector of covariates for respondent $i$ that we hypothesize might moderate the treatment effect. In our application, $\bm{X}_{i}$ included political party, education, demographics of their ZIP code, ethnicity, and Hispanic prejudice score (see Section 5.1 for details).
The randomness in our data, $(Y_{i},\bm{T}_{i},\bm{X}_{i})$ comes from two sources: random sampling of units into the study and random assignment of units to treatments. For simplicity, we assume units are sampled via simple random sampling (though our method can incorporate sampling weights). The randomization of treatment assignment implies $\{Y_{i}(\bm{t})\}_{\bm{t}\in\mathcal{T}}\ \mbox{$\perp\!\!\!\perp$}\ \bm{T}_{i}$ for each $i$ where the exact mode of randomization will determine the distribution of $\bm{T}_{i}$. In many conjoint experiments, researchers independently and uniformly randomize each factor. However, in some cases including our application, researchers may exclude certain unrealistic combinations of factor levels (e.g., doctor without a college degree), leading to the dependence between factors. In all cases, researchers have complete knowledge of the randomization distribution of the factorial treatment variables.
Based on random sampling and random treatment assignment alone, we can conduct valid inference for marginal treatment effects of interest using simple regression or difference-in-means estimator (see Hainmueller, Hopkins and Yamamoto, 2014). If we wish to explore treatment effect heterogeneity across treatments and covariates, however, a model-based approach is useful. We next introduce our model, which will allow us to explore heterogeneous effects in a principled manner while also handling the high-dimensional nature of the data.
3.2 General Framework
The most basic causal quantity of interest is the AMCE, which is defined for any given factor $j$ as
$\delta_{j}(l,l^{\prime})\ =\ \mathbb{E}[Y_{i}(T_{ij}=l,\bm{T}_{i,-j})-Y_{i}(T_{ij}=l^{\prime},\bm{T}_{i,-j})],$ (3.1)
where $l\neq l^{\prime}\in\mathcal{T}_{j}$ with $\mathcal{T}_{j}$ representing the support of the randomization distribution for $T_{j}$. The expectation in Equation (3.1) is taken over the distribution of the other factors $\bm{T}_{i,-j}$ as well as the random sampling of units from the population. Thus, the AMCE averages over two sources of causal heterogeneity—heterogeneity across treatment combinations and across units. Different treatment combinations may have distinct impacts on units with varying characteristics. Our goal is to model these potentially complex heterogeneous treatment effects using an interpretable model.
We propose to model heterogeneous treatment effects based on $K$ distinct treatment effect patterns where $K\geq 2$ is chosen by a researcher, based on their desired granularity of heterogeneity. This approach, which is based on a fixed number of subgroups to characterize treatment effect heterogeneity, is commonly used by empirical researchers. Others have studied various methodological aspects of this approach albeit in the context of binary treatment (Chernozhukov et al., 2019; Imai and Li, 2025).
Our goal is to summarize the treatment effect heterogeneity by dividing the population into $K$ subpopulations and characterizing these groups based on a set of pre-treatment covariates or “moderators” denoted by $\bm{X}_{i}$. In particular, we would like to construct $K$ groups such that across-group treatment effect heterogeneity is maximized while minimizing the within-group heterogeneity. Since the treatments of interest are high-dimensional, we focus on finding maximally heterogeneous groups in terms of average potential outcomes rather than their contrasts. We can then estimate any treatment effects of interest within each group.
Let $Z_{i}\in\{1,\cdots,K\}$ denote the latent group membership of unit $i$ and $\mathcal{Z}=\{Z_{i}\}_{i=1}^{n}$. We use $\zeta_{k}(\bm{t})=\mathbb{E}[Y_{i}(\bm{t})\mid Z_{i}=k]$ to represent the average potential outcome under treatment $\bm{t}$ for group $k$. Under the randomization of $T_{i}$, define the estimated within-group average outcome under treatment $t$ for group $k$ and the estimated overall average outcome as $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})=\sum_{i=1}^{N}I\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}Y_{i}/\sum_{i=1}^{N}I\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}$ and $\widehat{\overline{Y}}(\bm{t})=\sum_{i=1}^{N}I\{\bm{T}_{i}=\bm{t}\}Y_{i}/\sum_{i=1}^{N}I\{\bm{T}_{i}=\bm{t}\}$, respectively.
Given the number of groups $K$ selected by researchers, we show how to find maximally heterogeneous groups in terms of potential outcomes. The following proposition establishes that maximizing the Kullback-Leibler (KL) divergence of potential outcomes between groups is equivalent to maximizing the log-likelihood over groups and their centroids. We emphasize that this equivalence result does not assume the existence of a “correct” number of groups.
Proposition 1 (Finding maximally heterogeneous groups).
Maximally heterogeneous groups in the terms of the KL divergence of potential outcomes can be found by maximizing the log-likelihood function over the group membership and the centroids of groups,
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)\right\}$ (3.2)
$=$ $\operatornamewithlimits{argmax}_{\mathcal{Z}}\sum_{k=1}^{K}\sup_{\zeta_{k}}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\left[Y_{i}\log\zeta_{k}(\bm{T}_{i})+(1-Y_{i})\log\{1-\zeta_{k}(\bm{T}_{i})\}\right]$
where $Y_{i}$ is binary, the KL divergence of two Bernoulli distributions with means $\mu_{1}$ and $\mu_{2}$ is given by $\mathrm{KL}\left(\mu_{1}\|\mu_{2}\right)\ =\ \mu_{1}\log\mu_{1}/\mu_{2}+(1-\mu_{1})\log(1-\mu_{1})/(1-\mu_{2})$, and $\{\hat{\zeta}_{k}(\bm{t};\mathcal{Z})\}_{k=1}^{K}$ denotes the maximizers of the right hand side of Equation (3.2) given $\mathcal{Z}$.
Section C of the Supplementary Material provides a proof of a more general result for the natural exponential family distributions (see Chi, Chi and Baraniuk, 2016, for a similar result in the Gaussian case). The log-likelihood formulation is equivalent to the classification maximum likelihood approach in mixture modeling (McLachlan, 1982).
We now extend the above equivalence result to the settings in which we further model the group membership $Z_{i}$ using a set of moderators $\bm{X}_{i}$, i.e., $\pi_{k}(\bm{x})=\mathrm{Pr}(Z_{i}=k\mid\bm{X}_{i}=\bm{x})$ for $k=1,2,\ldots,K$. Such a model helps characterize and understand the types of units that comprise each group. The next proposition shows that maximizing the log-likelihood function of this extended model is equivalent to finding $K$ maximally heterogeneous groups such that the group memberships are predicted well by the moderators.
Proposition 2 (Finding maximally heterogeneous groups with moderators).
Suppose that we extend the setting of Proposition 1 and additionally model the conditional probability of each individual’s group membership given categorical moderators $\{\pi_{k}(\bm{X}_{i})\}_{k=1}^{K}$. Then, maximally heterogeneous groups in terms of the KL divergence of potential outcomes with the entropy of group membership probabilities as a penalty term can be found by maximizing the log-likelihood function of the extended model,
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)-\sum_{i=1}^{N}H(\{\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z})\}_{k=1}^{K})\right\}$ (3.3)
$=$ $\operatornamewithlimits{argmax}_{\mathcal{Z}}\sum_{k=1}^{K}\sup_{\zeta_{k},\pi_{k}}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\left[Y_{i}\log\zeta_{k}(\bm{T}_{i})+(1-Y_{i})\log\{1-\zeta_{k}(\bm{T}_{i})\}+\log\pi_{k}(\bm{X}_{i})\right]$
where $H(\{p_{k}\}_{k=1}^{K})=-\sum_{k=1}^{K}p_{k}\log p_{k}$ (by convention, if $p_{k}=0$, then $p_{k}\log p_{k}=0$) is the entropy, and $\hat{\pi}_{k}(\bm{x};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{X}_{i}=\bm{x}\}/\sum_{i=1}^{N}\bm{1}\{\bm{X}_{i}=\bm{x}\}$ and $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})$ are the maximizers of the log-likelihood function of the right hand side of Equation (3.3) given $\mathcal{Z}$.
Proof is given in Section D of the Supplementary Material. Since the entropy $H(\{\hat{\pi}_{k}(\bm{x};\mathcal{Z})\}_{k=1}^{K})$ is maximized when $\hat{\pi}_{k}(\bm{x})=1/K$, Proposition 2 shows that adding a group membership model based on moderators encourages finding groups whose memberships are well predicted by the moderators.
Direct optimization of Equations (3.2) and (3.3) over $\mathcal{Z}$ has been studied under the name of “classification maximum likelihood” in the literature on mixture models (McLachlan, 1982). For completeness, Section G.3 of the Supplementary Material provides an estimation algorithm for this approach, which modifies the proposed algorithm described in Section 3.5. Unfortunately, the classification maximum likelihood approach suffers from the incidental parameter problem because the cardinality of $\mathcal{Z}$ increases with the sample size $N$, leading to an asymptotic bias and inconsistency (Bryant and Williamson, 1978).
To address this problem, a dominant approach in the literature is Bayesian, treating the right hand side of Equation (3.3) as a log-posterior that consists of a log-likelihood and a log-prior over $\mathcal{Z}$, i.e., $\mathrm{Pr}(Z_{i}=k\mid\bm{X}_{i})=\pi_{k}(\bm{X}_{i})$. By marginalizing out $\mathcal{Z}$, we avoid the incidental parameter problem, yielding the objective function known as a mixture maximum likelihood (McLachlan, 1982).
The model is called “mixture-of-experts” when $\pi_{k}$ depends on $\bm{X}_{i}$ (Gormley and Frühwirth-Schnatter, 2019) with the following objective function,
$\{\hat{\zeta}_{k},\hat{\pi}_{k}\}_{k=1}^{K}=\operatornamewithlimits{argmax}_{\{\zeta_{k},\pi_{k}\}_{k=1}^{K}}\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right].$ (3.4)
While this setup no longer appears to provide a direct characterization of the optimal groups, Proposition 3 shows that a mixture-of-experts model finds maximally heterogeneous groups as in Proposition 2 but with an additional penalty that encourages less extreme posterior probabilities of group memberships.
Proposition 3 (Finding maximally heterogeneous groups with a mixture of experts).
Maximizing the likelihood function under a mixture-of-experts model is equivalent to finding maximally heterogeneous groups as in Proposition 2 with an additional penalty. That is, the following equality holds for any $\mathcal{Z}$,
$\operatornamewithlimits{argmax}_{\zeta,\pi}\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right]$
$=$ $\operatornamewithlimits{argmax}_{\zeta,\pi}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbf{1}\{Z_{i}=k\}\left[Y_{i}\log\zeta_{k}(\bm{T}_{i})+(1-Y_{i})\log\{1-\zeta_{k}(\bm{T}_{i})\}\right.$
$\left.\hskip 72.26999pt+\log\pi_{k}(\bm{X}_{i})-\log\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\{\zeta_{k^{\prime}},\pi_{k^{\prime}}\}_{k^{\prime}=1}^{K})\right],$
where
$\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\{\zeta_{k^{\prime}},\pi_{k^{\prime}}\}_{k^{\prime}=1}^{K})\ =$ $\mathrm{Pr}(Z_{i}=k\mid Y_{i},\bm{T}_{i},\bm{X}_{i},\{\zeta_{k^{\prime}},\pi_{k^{\prime}}\}_{k^{\prime}=1}^{K})$
[…]
is the posterior membership probability for group $k$.
Proof of the proposition directly follows from a well-known identity (e.g., Celeux, Frühwirth-Schnatter and Robert, 2019), and hence is omitted. The equality in Proposition 3 holds for any group membership $\mathcal{Z}$, including its maximum-a-posteriori (MAP) estimate, i.e.,
$\hat{Z}_{i}=\operatornamewithlimits{argmax}_{k}\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\{\hat{\zeta}_{k^{\prime}},\hat{\pi}_{k^{\prime}}\}_{k^{\prime}=1}^{K})$. Thus, our proposed model can be seen as finding maximally heterogeneous groups while imposing a penalty that encourages finding groups that are well predicted by the moderators $\bm{X}_{i}$ but with less extreme group membership probabilities based on the data.
All together, our results provide a justification for using a mixture-of-experts model for heterogeneous effect estimation under the settings with high-dimensional treatments. We emphasize that a primary motivation for the use of Bayesian approach is to resolve the incidental parameter problem with classification maximum likelihood. Importantly, the results above do not assume a specific data generating process. Instead, we have shown that given the number of groups and appropriate prior distributions, researchers can find maximally heterogeneous groups by fitting a mixture-of-experts model.
3.3 Model Specification
Since $\mathcal{T}$ is high-dimensional, many treatment combinations are unobserved with a typical sample size. Thus, nonparametric estimation is not applicable. We, therefore, model $\zeta_{k}(\bm{t})$ using a regularized logistic regression where an ANOVA-style sum-to-zero constraint is imposed separately for each factor to facilitate merging of different levels within each factor. This modeling strategy identifies a relatively small number of treatment combinations while avoiding the specification of a baseline level for each factor (Egami and Imai, 2019). The interpretation of $\zeta_{k}(\bm{t})$ under this model is still the average of potential outcome under treatment $\bm{t}$ in group $k$. Note that we do not assume homogeneity of outcomes or effects within each group.
We use a multinomial logistic regression for $\pi_{k}(\bm{x})$:
$\zeta_{k}(\bm{T}_{i})\ =\ \frac{\exp(\psi_{k}(\bm{T}_{i}))}{1+\exp(\psi_{k}(\bm{T}_{i}))},\quad\text{and}\quad\pi_{k}(\bm{X}_{i})=\frac{\exp(\bm{X}_{i}^{\top}\bm{\phi}_{k})}{\sum_{k^{\prime}=1}^{K}\exp(\bm{X}_{i}^{\top}\bm{\phi}_{k^{\prime}})},$ (3.5)
where $\bm{\phi}_{1}=\bm{0}$ for identification. For $\psi_{k}(\bm{T}_{i})$, we assume an additive model and include both main effects and two-way interaction effects with a common intercept $\mu$ shared across all groups,
$\psi_{k}(\bm{T}_{i})$ $\ =\ \mu+\sum_{j=1}^{J}\sum_{l=0}^{L_{j}-1}\mathbf{1}\{T_{ij}=l\}\beta^{j}_{kl}+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\sum_{l=0}^{L_{j}-1}\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}\beta^{jj^{\prime}}_{kll^{\prime}}$
$\ =\ \mu+\tilde{\bm{T}}_{i}^{\top}\bm{\beta}_{k},$
for each $k=1,2,\ldots,K$ where $\tilde{\bm{T}}_{i}$ is the vector of indicators, $\mathbf{1}\{T_{ij}=l\}$ and $\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}$, and $\bm{\beta}_{k}$ is a stacked column vector containing all coefficients for group $k$. Inclusion of higher-order interactions is straightforward (see Section E of the Supplementary Material) and hence is omitted in the main paper for notational simplicity.
For identification, we use the following ANOVA-type sum-to-zero constraints,
$\sum_{l=0}^{L_{j}-1}\beta^{j}_{kl}=0,\quad\text{and}\quad\sum_{l=0}^{L_{j}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=0,$ (3.6)
for $j,j^{\prime}=1,2,\ldots,J$ with $j^{\prime}>j$. We write them compactly as,
[…]
where each row of $\bm{C}^{\top}\bm{\beta}_{k}$ corresponds to one of the constraints given in Equation (3.6).
3.4 Sparsity-inducing Prior
Given the high dimensionality of this model, we use a sparsity-inducing prior. In our application, we have a total of 315 $\beta$ coefficients for each group. In factorial experiments, it is desirable to regularize the model such that certain levels of each factor are fused together when their main effects and all interactions are similar (Post and Bondell, 2013; Egami and Imai, 2019). For example, we would like to fuse levels $l_{1}$ and $l_{2}$ of factor $j$ if $\beta^{j}_{l_{1}}\approx\beta^{j}_{l_{2}}$ and $\beta^{jj^{\prime}}_{l_{1}l^{\prime}}\approx\beta^{jj^{\prime}}_{l_{2}l^{\prime}}$for all other factors $j^{\prime}$ and all of its levels $l^{\prime}$.
We encourage such fusion by applying a structured sparsity approach of Goplerud (2021) that generalizes the group and overlapping group LASSO (e.g., Yuan and Lin, 2006; Yan and Bien, 2017) while allowing positive semi-definite penalty matrices. For computational tractability, we use $\ell_{2}$ norm instead of the $\ell_{\infty}$ norm, which is used in GASH-ANOVA (Post and Bondell, 2013). An additional benefit of the use of regularization is that it gives us some protection against finding spurious relations (see Gelman, Hill and Yajima, 2012).
For illustration, consider a simple example with one group and two factors—factor one has three levels and factor two has two levels. In this case, our penalty contains four terms,
$\begin{split}&\sqrt{(\beta^{1}_{0}-\beta^{1}_{1})^{2}+(\beta^{12}_{00}-\beta^{12}_{10})^{2}+(\beta^{12}_{01}-\beta^{12}_{11})^{2}}\\ ~{}+~{}&\sqrt{(\beta^{1}_{0}-\beta^{1}_{2})^{2}+(\beta^{12}_{00}-\beta^{12}_{20})^{2}+(\beta^{12}_{01}-\beta^{12}_{21})^{2}}\\ ~{}+~{}&\sqrt{(\beta^{1}_{1}-\beta^{1}_{2})^{2}+(\beta^{12}_{10}-\beta^{12}_{20})^{2}+(\beta^{12}_{11}-\beta^{12}_{21})^{2}}\\ ~{}+~{}&\sqrt{(\beta^{2}_{0}-\beta^{2}_{1})^{2}+(\beta^{12}_{00}-\beta^{12}_{01})^{2}+(\beta^{12}_{10}-\beta^{12}_{11})^{2}+(\beta^{12}_{20}-\beta^{12}_{21})^{2}}.\end{split}$
The first three terms encourages the pairwise fusion of the levels of factor one whereas the fourth encourages the fusion of the two levels of factor two. For compact notation, the penalty can also be written using the sum of Euclidean norms of quadratic forms,
$||\bm{\beta}^{\top}\bm{F}_{1}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{2}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{3}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{4}\bm{\beta}||_{2},$
where $\bm{F}_{1},\bm{F}_{2},\bm{F}_{3}$ are appropriate positive semi-definite matrices to encourage the fusion of the pairs of levels in factor one and $\bm{F}_{4}$ encourages the fusion of the two levels in factor two, and
[…]
We generalize this formulation to an arbitrary number of factors and factor levels. For each factor that contains $L_{j}$ levels, we have $\binom{L_{j}}{2}$ penalty matrices to encourage pairwise fusion. Imposing additional constraints is a simple extension; for example, for ordered factors, one might use penalties that penalize the differences between adjacent levels (e.g. $l$ and $l+1$). Let $G=\sum_{j=1}^{J}\binom{L_{j}}{2}$ represent the total number of penalty matrices. For $g=1,2,\ldots,G$, we use $\bm{F}_{g}$ to denote a penalty matrix such that $\sqrt{\bm{\beta}^{\top}\bm{F}_{g}\bm{\beta}}$ is equivalent to the $\ell_{2}$ norm on the vector of differences between all main effects and interactions containing a main effect. We note that $\{\bm{F}_{g}\}_{g=1}^{G}$ is not directly chosen but rather are determined by factors in the experiment ($J$, $L_{j}$, whether $j$ is ordered or unordered) and the included interactions (as well as the use of “latent overlapping groups”; see Section H.4 of the Supplementary Material.
We interpret this penalty as a prior under our Bayesian framework described in Section 3.2,
$p\left(\bm{\beta}_{k}\mid\{\bm{\phi}_{k}\}_{k=2}^{K}\right)\ \propto\ \left(\lambda{\bar{\pi}_{k}^{\gamma}}\right)^{m}\exp\left(-\lambda\bar{\pi}_{k}^{\gamma}\sum_{g=1}^{G}\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}\right),$ (3.8)
where $\bar{\pi}_{k}=\sum_{i=1}^{N}\pi_{k}(\bm{X}_{i})/N$ and $m=\text{rank}\left([\bm{F}_{1},\cdots,\bm{F}_{G}]\right)$. We follow existing work in allowing the penalty on the treatment effects $\bm{\beta}_{k}$ to be scaled by the group-membership size $\bar{\pi}_{k}$ when $\gamma=1$ (Khalili and Chen, 2007; Städler, Bühlmann and Van De Geer, 2010). On the other hand, when $\gamma=0$ the $\bar{\pi}_{k}$ disappears, implying no use of the $\bm{X}_{i}$ in the prior. We note that the prior on $p(\bm{\beta}\mid\{\bm{\phi}_{k}\}_{k=2}^{K})$ is guaranteed to be proper when all pairwise fusions are encouraged by $\{\bm{F}_{g}\}_{g=1}^{G}$, although in other circumstances it may be improper (Goplerud, 2021). Section F of the Supplementary Material provides additional details. Following Zahid and Tutz (2013), we use a normal prior distribution for the coefficients for the moderators.
The resulting regularization is invariant to the choice of baseline group $\bm{\phi}_{1}=\bm{0}$, which is the first row of the $K\times p_{x}$ coefficient matrix $\bm{\phi}$. The prior distribution is given by
$p(\{\bm{\phi}_{k}\}_{k=2}^{K})\ \propto\ \exp\left(-\frac{\sigma^{2}_{\phi}}{2}\sum_{l=1}^{p_{x}}[\bm{\phi}_{2l},\cdots,\bm{\phi}_{Kl}]^{\top}\bm{\Sigma}_{\phi}[\bm{\phi}_{2l},\cdots,\bm{\phi}_{Kl}]\right),$ (3.9)
where $\bm{\Sigma}_{\phi}$ is a $(K-1)\times(K-1)$ matrix with $[\bm{\Sigma}_{\phi}]_{kk^{\prime}}=(K-1)/K$ if $k=k^{\prime}$ and $[\bm{\Sigma}_{\phi}]_{kk^{\prime}}=-1/K$ otherwise. We set $\sigma^{2}_{\phi}$ to $1/4$ for a relatively diffuse prior.
As noted in a recent survey, “ensuring generic identifiability for general [mixture of expert] models remains a challenging issue” (Gormley and Frühwirth-Schnatter, 2019, p. 294). Although mixtures with a Bernoulli outcome variable are generally unidentifiable, several aspects of our methodology are expected to alleviate the identifiability problem. First, a typical conjoint analysis has repeated observations per unit $i$ (Grün and Leisch, 2008). Second, our model is a mixture of experts rather than a mixture model (Jiang and Tanner, 1999). Third, our treatment variables, which act as covariates in a mixture of experts, are randomized and hence uncorrelated with one another. Lastly, our model regularizes the coefficients through an informative prior. While a formal identifiability analysis of our model is beyond the scope of this paper, the simulation analysis (Section 4) shows that our model can accurately recover the coefficients in a realistic setting. It is also possible to use a bootstrap-based procedure to examine the identifiability issue in a specific setting (Grün and Leisch, 2008).
3.5 Estimation and Inference
We fit our model by finding a maximum of the log-posterior using an extension of the Expectation-Maximization (EM; Dempster, Laird and Rubin 1977) algorithm known as the Alternating Expectation-Conditional Maximization (AECM; Meng and van Dyk 1997) algorithm. Equation (3.10) defines our (observed) log-posterior using the terms defined in Equations (3.3), (3.8), and (3.9), where we collect all model parameters as $\bm{\theta}$:
$\begin{split}\log p\left(\bm{\theta}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)\ =\ &\sum_{i=1}^{N}\log\left[\sum_{k=1}^{K}\pi_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right]+\\ &\sum_{k=1}^{K}\log p(\bm{\beta}_{k}\mid\{\bm{\phi}\}_{k=2}^{K})+\log p(\{\bm{\phi}_{k}\}_{k=2}^{K})+\text{const}.\end{split}$ (3.10)
For now, we assume the value of regularization parameter $\lambda$ is fixed, although we discuss this issue in Section 3.6. The linear constraints on $\bm{\beta}_{k}$ given in Equation (3.7) still hold but are suppressed for notational simplicity.
Section G of the Supplementary Material provides a full derivation of our AECM algorithm; each iteration involves two cycles where the data augmentation scheme enables iterative updating of the treatment effect parameters $\bm{\beta}$ and moderators $\bm{\phi}$. After augmenting with missing data, the update for $\bm{\beta}$ can be done using ridge regression; Section G.1 addresses the linear constraints imposed by $\bm{C}^{T}\bm{\beta}_{k}=\bm{0}$. The update for $\bm{\phi}$ can be performed using a modified version of a multinomial logistic regression based on a standard optimizer (e.g., L-BFGS) (see Section G.2).
3.6 Additional Considerations
Since fitting the proposed model is computationally expensive, we use the Bayesian Information Criteria (BIC), rather than cross validation, to select the value of the regularization parameter $\lambda$ (Khalili and Chen, 2007; Khalili, 2010; Chamroukhi and Huynh, 2019). Section G.4 of the Supplementary Material presents our degrees-of-freedom estimator and explains how we tune $\lambda$ using Bayesian model-based optimization. Section G.5 discusses additional details of our EM algorithm including initialization and techniques to accelerate convergence.
We extend the above model and estimation algorithm to accommodate common features of conjoint analysis: (1) repeated observations for each individual respondent (Section H.1 of the Supplementary Material), (2) a forced choice conjoint design (Section H.2), and (3) standardization weights for factors with different numbers of levels $L_{j}$ (Section H.3). Lastly, our experience suggests that the proposed penalty function, which consists of overlapping groups, often finds highly sparse solutions. Section H.4 details the integration of the latent overlapping group formulation of Yan and Bien (2017) into our framework to address this issue.
Once the model parameters are estimated, we can compute quantities of interest such as the AMCEs, defined in Equation (3.1). We do this separately for each group, such that $\delta_{jk}(l,l^{\prime})$ is the AMCE for factor $j$, changing from level $l^{\prime}$ to $l$ in group $k$. Our estimator is the average of the estimated difference in predicted responses when changing from level $l^{\prime}$ to $l$ of factor $j$, where the average is taken over the empirical distribution of the assignment on the other factors. This estimation is described in more detail in Section I of the Supplementary Material under various settings. We can use the empirical distribution here because treatment is randomly assigned.
To quantify the uncertainty of the parameter estimates, we rely on a quadratic approximation to the log posterior distribution. To ensure its differentiability, we follow a standard approach in the regularized regression literature (e.g., Fan and Li, 2001) and fuse pairwise factor levels that are sufficiently close together. Section J of the Supplementary Material describes this process, deriving the Hessian of the log-posterior using Louis (1982)’s method and then using the delta method for inference on other of quantities of interest, e.g., the AMCE.
Finally, in principle, our framework does not assume a “correct” data generating process. The choice of number of groups $K$ should depend on the desired granularity of discovered heterogeneity, with more groups leading to finer levels of heterogeneity. Similarly, the choice of moderators should reflect the researcher’s substantive interests. Section K.3 of the Supplementary Material shows performance of our method across different values of $K$ and different specifications of the moderators when the true data generating process is a mixture model. As expected, the bias of AMCE is not affected by changing the specification of these parameters. However, there are some impacts on the estimation of conditional effects in terms of precision.
Common data-driven approaches for choosing $K$ include use of an information criterion such as the BIC; however, while we find that these approaches work well under simulation settings (see Section K.3.1 for demonstration), they can perform poorly in practice (see Section L), especially when the component densities are mis-specified or not especially well separated (Celeux, Frühwirth-Schnatter and Robert, 2019). Thus, even if a data-driven heuristic is used as a guide for choosing $K$, we suggest comparing different $K$ as illustrated in Section 5.
4 Simulations
We explore the performance of our method using a simple but realistic simulation study. Specifically, we consider the case of a conjoint experiment with ten factors ($J=10$) each with three levels ($L_{j}=3$). To evaluate the performance of the proposed method, we consider two different settings; in the first, we assume there are 1,000 respondents who each perform five comparison tasks. In the second, we assume a larger experiment with 2,000 respondents who each perform ten tasks.
In all cases, we assume that the data generating process follows a mixture of experts model with three groups ($K=3$). We calibrate the true $\bm{\beta}_{k}$ such that the implied average marginal component effects (AMCE) are comparable in magnitude to the empirical effects presented in Section 5. We use a set of five correlated continuous moderators and an intercept to again mimic a realistic empirical setting and choose $\{\bm{\phi}_{k}\}_{k=2}^{3}$ to relatively clearly separate respondents into different groups. Section K of the Supplementary Material presents complete description of the simulation settings and the true parameter values used for the $\bm{\beta}_{k}$ and marginal effects.
For each sample size, we independently generate 1,000 simulated data sets by drawing $N$ observations of moderators, randomly assigning a group membership to each observation based on the implied probabilities given their moderators, and generating the observed treatment profiles completely at random. We fit our model to the data with $K=3$ and examine the average marginal component effects in each group with respect to the first baseline level.
Figure 2: The empirical performance of the proposed estimator on simulated data. The black squares indicate the effects estimated for each group with the smaller sample size (1,000 people completing 5 tasks); the red crosses indicate effects estimated with the larger sample size (2,000 people completing 10 tasks).
Fig 2 summarizes our results (see Section K.2 for the results regarding the estimated coefficients $\bm{\beta}_{k}$). The left panel illustrates a high correlation between the estimated effects and their true values ($\rho=0.995$ for smaller sample size; $\rho=0.999$ for larger sample size). While the performance overall is reasonably strong, we see that even when the dataset is large there is some degree of attenuation bias due to shrinkage.
The right panel shows the frequentist evaluation of our Bayesian posterior standard deviations. We compare the average posterior standard deviation against the standard deviation of the estimated effects across the 1,000 Monte Carlo simulations. The average posterior standard deviations are noticeably smaller than the standard deviation of the estimates when the sample size is small. For the large sample size, however, our approximate Bayesian posterior standard deviations in this simulated example are roughly the same magnitude of the standard deviation of the sampling distribution of the estimator.
Even though our method’s frequentist coverage is somewhat below the nominal level in small samples, this undercoverage appears to be primarily attributable to the shrinkage bias in our regularized estimation rather than the large sample discrepancy between our posterior standard deviations and the corresponding standard deviation of sampling distribution.
Section K.2 explores one way to address the limitations of the default estimator by exploring sample splitting and refitting the model given the estimated sparsity pattern (i.e., which levels are fused together) and moderator effects ($\{\bm{\phi}_{k}\}_{k=2}^{K}$) on half of the data. This results in smaller bias and improved coverage at both sample sizes.
Section K.3 explores how when the true data generating process is a mixture model, the “wrong” choice of $K$, e.g. $K\in\{1,2,4\}$, as well as not using moderators (i.e., $\bm{X}_{i}=1$) or using moderators in a different specification than the true model impacts our results. In both settings, there is limited impact in terms of bias in terms of estimating the AMCE, although both types of misspecification incur a penalty in terms of root mean-squared error.
5 Empirical Analysis
In this section, we apply our methodology to the immigration conjoint data introduced in Section 2. We find evidence of effect heterogeneity for immigrant choice based on respondent characteristics. In particular, the immigrant’s country of origin plays a greater role in forming the immigration preference of respondents with increased prejudice, as measured by a Hispanic prejudice score. Outside of this group, which accounts for about one third of the respondents, the country of origin factor plays a much smaller role.
5.1 Data and Model
Following the original analysis, our model includes indicator variables for each factor and interactions between country and reason of application factors as well as those between education and job factors in order to account for the restricted randomization. We additionally include interactions between country and job as well as those between country and education, in accordance with the skill premium theory of Newman and Malhotra (2019). This theory hypothesizes that prejudiced individuals prefer highly skilled immigrants only for certain immigrant countries. This results in a total of 41 AMCEs and 222 average marginal interaction effects (AMIEs) for each group.
For modeling group membership, we include the respondents’ political party, education, demographics of their ZIP code (we follow the original analysis and include the variables indicating whether respondents’ ZIP code had few immigrants, meaning $<5\%$, and for those from ZIPs with more than $5\%$ foreign-born, whether the majority were from Latin America), ethnicity, and Hispanic prejudice score. The Hispanic prejudice score was used by Newman and Malhotra (2019), though we negate it to make lower values correspond to lower prejudice for easier interpretation. The score is based on a standardized (and negated) feeling thermometer for Hispanics. The score ranges from $-1.61$ to $2.11$ for our sample, where higher scores indicate higher levels of prejudice.
We remove respondents who are themselves Hispanic since the Hispanic prejudice score was not measured for these respondents. After removing entries with missing data, we have a sample of 1,069 respondents. Most respondents evaluated five pairs of profiles, though five respondents have fewer responses in the data set used. The total number of observations is 5,337 pairs of profiles. We do not incorporate the survey weights into our analysis to better demonstrate our methods though it is possible to include them.
The original experiment was conducted using the forced choice design, in which a respondent chooses one profile out of a pair of immigrant profiles. We follow Egami and Imai (2019) and model the choice as a function of differences in treatments as follows,
$\psi_{k}(\bm{T}_{i}^{L},\bm{T}_{i}^{R})\ =$ $\mu+\sum_{j=1}^{J}\sum_{l\in L_{j}}\beta^{j}_{kl}\left(\mathbf{1}\left\{T^{L}_{ij}=l\right\}-\mathbf{1}\left\{T^{R}_{ij}=l\right\}\right)$
$+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\mathbf{1}\left\{\mathcal{I}(j,j^{\prime})\right\}\sum_{l\in L_{j}}\sum_{l^{\prime}\in L_{j^{\prime}}}\beta^{jj^{\prime}}_{kll^{\prime}}\left(\mathbf{1}\left\{T^{L}_{ij}=l,T^{L}_{ij^{\prime}}=l^{\prime}\right\}-\mathbf{1}\left\{T^{R}_{ij}=l,T^{R}_{ij^{\prime}}=l^{\prime}\right\}\right),$
where $\bm{T}_{i}^{L}$ and $\bm{T}_{i}^{R}$ represent the factors for the left and right profiles and $\mathcal{I}(j,j^{\prime})=1$ if an interaction between $j$ and $j^{\prime}$ is include in the model. The outcome variable $Y_{i}$ is equal to 1 if the left profile is selected and is equal to 0 if the right profile is chosen.
To account for randomization restrictions, we include interactions between country of origin and reason for applying as well as between job and education. To test relevant theories, we include additional interactions between country of origin and job as well as country of origin and education. These interaction effects proved to be very small in magnitude (see Section L of the Supplementary Material). Thus, we do not explore higher order interactions given the commonly adopted principle of hierarchy and sparsity (Wu and Hamada, 2021), which implies that lower-order effects are expected to be more significant than higher-order effects and we should expect an even smaller number of nonzero effects. With this linear predictor formulation, the estimation and inference proceed as explained in Section 3.
We conduct two analyses, one with two groups and the other with three groups. These two models perform equally well in terms of out-of-sample classification, a data-driven measure that can be used to choose the number of groups. Using more than three groups does not give improved substantive insights and provides little improvement in model performance. As noted previously, each analysis optimizes the BIC to calibrate the amount of regularization and employs standardization weights to account for factors with different number of levels (see Sections G.4 and H.3 of the Supplementary Material, respectively, for details). We treat education and job experience as ordered factors and only penalize the differences between adjacent levels.
We report our findings using only the full data estimates, i.e., without the sample splitting explored in Section K.2. Initial experiments found that the results were somewhat sensitive to specific folds chosen, and thus we report only the full data results in the main text. Section L illustrates the distribution of estimates across twenty different sample splits.
5.2 Estimated Heterogeneity
We focus on the AMCE for each factor as the quantity of interest and separately estimate it for each group. Under our model for the forced choice design, the AMCE of level $l$ versus level $l^{\prime}$ of factor $j$ within group $k$ can be written as,
$\delta_{jk}(l,l^{\prime})$ $\ =\ \frac{1}{2}\mathbb{E}\left[\left\{\Pr\left(Y_{i}=1\mid Z_{i}=k,T^{L}_{ij}=l,\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i}\right)-\Pr\left(Y_{i}=1\mid Z_{i}=k,T^{L}_{ij}=l^{\prime},\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i}\right)\right\}\right.$
+$\hskip 36.135pt\left.+\left\{\Pr\left(Y_{i}=0\mid Z_{i}=k,T^{R}_{ij}=l,\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i}\right)-\Pr\left(Y_{i}=0\mid Z_{i}=k,T^{R}_{ij}=l^{\prime},\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i}\right)\right\}\right].$
The expectation is over the population of respondents and the distribution of the factors not involved in this AMCE. That is, we compute the AMCE separately for the left and right profiles and then average them to obtain the overall AMCE. We estimate this quantity using the fitted model and averaging over the empirical distribution of the factorial treatments.
Figure 3: Estimated average marginal component effects using a two-group (left) and three-group (right) analysis. The point estimates and 95% Bayesian credible intervals are shown. A solid circle represents the baseline level of each factor. Numbers after colons give average posterior predictive probabilities for each group.
Fig 3 presents the estimated AMCEs and their 95% Bayesian credible intervals for the two-group and three-group analyses in the left and right panels, respectively. Group 2 in the two-group analysis and Group 3 in the three-group analysis display stronger impacts of country of origin than the other groups. The respondents in these groups give the most preference to immigrants from Germany and the least preference to immigrants from Iraq (followed by Sudan). The significant negative effects of Iraq in Group 2 of the two-group analysis and Group 3 of the three-group analysis are consistent with the significant negative effect for Iraq found by Hainmueller and Hopkins (2015). The patterns we observe for the other factors are also similar for these two groups in the two analyses.
Across all groups, respondents prefer educated and experienced immigrants who already have contracts (over those who have no contracts or plans). Respondents also prefer immigrants who have better language skills, although this feature matters less for respondents in Group 1 of the three group analyses.
For both analyses, the respondents in Group 1 do not care much about immigrant’s country of origin. Instead, they place a greater emphasis on education and reason for immigration when compared to those in the other groups. While the differences between Groups 1 and 2 in the three-group analysis are generally substantively small, those in Group 2 appear to place more emphasis on education and prior entry without legal authorization. Those in Group 1, on the other hand, give a slight benefit to immigrants whose reason for immigration is persecution.
Figure 4: Ternary plot of the joint posterior predictive probability of belonging to each group in the three-group analysis (three axes) where the color of each dot represents the posterior predictive probability of belonging to Group 1 under the two-group analysis.
Indeed, for the three-group analysis, Groups 1 and 2 together correspond roughly to Group 1 of the two-group analysis. In fact, about 81% of the respondents who belong to Group 1 of the two-group analysis are the members of either Group 1 or 2 in the three-group analysis, using a weighted average of their estimated group membership posterior predictive probabilities.
Fig 4 visualizes these posterior predictive probabilities of group membership under the three-group analysis with each dot colored by the posterior predictive probability of belonging to Group 1 under the two-group analysis. According to this ternary plot, those observations that are likely to be part of Group 1 under the two-group analysis (i.e., red dots) are likely to be split between Groups 2 and 3 under the three group analysis. In contrast, those who have a high probability of belonging to Group 2 under the two-group analysis (i.e., blue dots) tend to be part of Group 3.
Fig 3 shows fusion of various factor levels due to regularization. The levels being fused appear sensible. For example, “doctor” and “research scientist,” both occupations requiring high levels of education, are consistently fused together. For education, use of the ordinal structure ensures only adjacent levels can be fused. We see sensible cut points for fusion; in the two group analysis, Group 1 differentiates individuals who have at least a college degree and Group 2 differentiates individuals who have at least a high school degree.
The comparison of AMCEs across subgroups can be misleading as they depend on the choice of baseline category (Leeper, Hobolt and Tilley, 2020). Section L of the Supplementary Material presents an alternative quantity that avoids issues of baseline dependency (marginal means; Leeper, Hobolt and Tilley 2020). The results are generally similar to AMCEs shown above.
Figure 5: Distribution of respondent characteristics for each group. Left set of plots shows weighted box plots of the Hispanic prejudice moderator within each group over the posterior predictive distribution using a two-group (left) and three-group (right) analysis. Right set of plots shows the distribution of categorical moderators within each group over the posterior predictive distribution using a two-group (left) and three-group (right) analysis.
5.3 Group Membership
Who belongs to each group? The left panel of Fig 5 shows the distribution of Hispanic prejudice score for each group weighted by the corresponding posterior predictive group membership probability for each individual respondent. The plot shows that for the two-group analysis, those with high prejudice score are more likely to be part of Group 2. For the three-group analyses, those with high prejudice are more likely to be in Group 3. This is consistent with the finding above that the respondents in those groups put more emphasis on immigrant’s country of origin.
The right panel of the figure shows the distribution of other respondent characteristics. In general, Group 2 in the two-group analysis and Group 3 in the three-group analysis consist of those who live in ZIP codes with few immigrants and have lower educational achievements. For the three-group analysis, those in Group 2 tend to be Republicans, whereas those in Group 1 are more likely to be Democrats. This is consistent with the finding of a larger penalty for entry without legal authorization in Group 2. Group 3 contains a mix of political ideologies, though it has more respondents who identify as Undecided/Independent/Other or not strong Republican than the other two groups.
Which respondent characteristics are predictive of the group membership? In addition to the covariate distribution for each group shown in Fig 5, we can also find how important each moderator is in predicting group membership, conditional on all other moderators. We examine how the predicted probabilities of group memberships change across respondents with different characteristics. Specifically, we estimate
$\mathbb{E}\left[\pi_{k}(X_{ij}=x_{1},\bm{X}_{i,-j})-\pi_{k}(X_{ij}=x_{0},\bm{X}_{i,-j})\right]$ (5.1)
where $x_{0}$ and $x_{1}$ are different values of covariate of interest $X_{ij}$. If $X_{ij}$ is a categorical variable, we set $x_{0}$ to the baseline level and $x_{1}$ to the level indicated on the vertical axis. If $X_{ij}$ is a continuous variable as in the case of the Hispanic prejudice score, then $x_{0}$ and $x_{1}$ represent the 25th and 75th percentile values. The solid arrows represent whether the corresponding 95% Bayesian credible interval covers zero or not. Section L of the Supplementary Material shows the effect of changing a moderator on the absolute value of the changes in predicted probabilities of group membership. In some cases, changing a moderator shows a small average change but a larger average of absolute changes.
Figure 6: The impact of moderator values on likelihood of being assigned to groups, for two-group (left two plots) and three-group (right three plots) analysis. Dark arrows indicate that there is a significant effect of the moderator on group membership, i.e., that the corresponding quantity defined in Equation (5.1) is statistically significant.
Consistent with the earlier findings, Fig 6 shows that those with high Hispanic prejudice scores tend to be part of Group 2 in the two-group analysis and Group 3 in the three-group analysis even after controlling for other moderators. These respondents are also less likely to be members of Group 1 in both analyses. Party ID also plays a statistically significant role (indicated by dark arrow). Controlling for other factors, in the three-group analysis, not strong Republicans tend to be part of Group 3 and more strong Democrats belonging to Group 1. On average, respondents in Group 1 have higher education in both analyses.
Finally, we estimate the average marginal interaction effects (AMIEs) between two factors (Egami and Imai, 2019), which can be computed by subtracting the two AMCEs from the average effect of changing the two factors of interest at the same time. Thus, the AMIE represents the additional effect of changing the two factors beyond the sum of the average effects of changing one of the factors alone. Formally, we can define the AMIE of changing factors $j$ and $j^{\prime}$ from levels $l_{j}$ and $l_{j^{\prime}}$ to levels $l_{j}^{\prime}$ and $l_{j^{\prime}}^{\prime}$, respectively, as follows,
$\mathbb{E}[Y_{i}(T_{ij}=l_{j},T_{ij^{\prime}}=l_{j^{\prime}},\bm{T}_{i,-j,-j^{\prime}})-Y_{i}(T_{ij}=l_{j}^{\prime},T_{ij^{\prime}}=l_{j^{\prime}}^{\prime},\bm{T}_{i,-j,-j^{\prime}})]-\delta_{j}(l_{j},l_{j}^{\prime})-\delta_{j^{\prime}}(l_{j^{\prime}},l_{j^{\prime}}^{\prime}).$
All of the AMIE effects found are quite small, so we do not present those results here. According to the skill-premium theory of Newman and Malhotra (2019), we expect to find an interaction between job and country or education and country, in at least some groups. Unfortunately, our analysis does not find support for this hypothesis.
5.4 Comparison to an alternative method
While there exist few methods to estimate heterogeneous effects of high-dimensional treatments, an exception is Robinson and Duch (2024), who develop a BART-based method for analyzing heterogeneity in conjoint experiments. The primary goal of their method is the estimation of the conditional average marginal effects (CAMCE) for each individual given their covariate values.
While our method is motivated by a different goal—finding an interpretable set of groups with distinctive treatment effects—our method can also produce estimates of the CAMCE for any set of covariates. The two methods can be compared in this task by examining CAMCE. Formally, under our model, the CAMCE for factor $j$ comparing levels $l$ and $l^{\prime}$ for covariates $\bm{X}_{i}$ is a weighted average of the group-specific AMCEs, denoted by $\delta_{jk}(l,l^{\prime})$.
$\mathrm{CAMCE}_{j}(l,l^{\prime};\bm{X}_{i})=\sum_{k=1}^{K}\delta_{jk}(l,l^{\prime})\pi_{k}(\bm{X}_{i})$ (5.2)
By plugging in our estimates $\hat{\pi}_{k}(\bm{X}_{i})$ and $\hat{\delta}_{jk}(l,l^{\prime})$, we can estimate the CAMCE.
Section B of the Supplementary Material compares the estimated CAMCE obtained from our method and Robinson and Duch (2024)’s (cjbart) using the same moderators and treatments. Our method discovers a considerable degree of heterogeneity in the CAMCEs whereas cjbart shows limited treatment effect variation for most countries. Under our model, the estimated heterogeneous effects are more strongly associated with predictors than cjbart; for example, our method finds a clear association, on average, between the estimated CAMCE and prejudice or party identification whereas cjbart does not.
6 Concluding Remarks
We have shown that a Bayesian mixture of regularized logistic regressions can be effectively used to estimate heterogeneous treatment effects of high-dimensional treatments. The proposed approach finds maximally heterogeneous groups and yields interpretable results, illuminating how different sets of treatments have heterogeneous impacts on distinct groups of units. We apply our methodology to conjoint analysis, which is a popular survey experiment. Our analysis shows that individuals with high prejudice score tend to discriminate against immigrants from certain non-European countries. These individuals tend to be less educated and live in areas with few immigrants. Future research should consider the derivation of optimal treatment rules in this setting as well as the empirical evaluation of such rules. Another important research agenda is the estimation of heterogeneous effects of high-dimensional treatments in observational studies.
Supplementary Material for
“Estimating Heterogeneous Causal Effects of High-Dimensional Treatments: Application to Conjoint Analysis”
Max Goplerud             Kosuke Imai             Nicole E. Pashley
Supplementary Material A The Details of the Immigration Conjoint Experiment
Attribute # of Levels Levels
Education 7 No formal education; Equivalent to completing fourth grade in the U.S.; Equivalent to completing eighth grade in the U.S.; Equivalent to completing high school in the U.S.; Equivalent to completing two years at college in the U.S.; Equivalent to completing a college degree in the U.S.; Equivalent to completing a graduate degree in the U.S.
Gender 2 Female; Male
Country of origin 10 Germany; France; Mexico; Philippines; Poland; India; China; Sudan; Somalia; Iraq
Language 4 During admission interview, this applicant spoke fluent English; During admission interview, this applicant spoke broken English; During admission interview, this applicant tried to speak English but was unable; During admission interview, this applicant spoke through an interpreter
Reason for Application 3 Reunite with family members already in U.S.; Seek better job in U.S.; Escape political/religious persecution
Profession 11 Gardener; Waiter; Nurse; Teacher; Child care provider; Janitor; Construction worker; Financial analyst; Research scientist; Doctor; Computer programmer
Job experience 4 No job training or prior experience; One to two years; Three to five years
Employment Plans 4 Has a contract with a U.S. employer; Does not have a contract with a U.S. employer, but has done job interviews; Will look for work after arriving in the U.S.; Has no plans to look for work at this time
Prior Trips to the U.S. 5 Never been to the U.S.; Entered the U.S. once before on a tourist visa; Entered the U.S. once before without legal authorization; Has visited the U.S. many times before on tourist visas; Spent six months with family members in the U.S.
Table A1: Table 1 in Hainmueller and Hopkins (2015). All attributes for immigrants and their levels.
Supplementary Material B Additional Results for Comparison with cjbart
We compare the performance of our method with that of Robinson and Duch (2024) whose method is implemented using an open-source software package, cjbart (Robinson and Duch, 2023). We use the same set of moderators and factors considered in our earlier analyses. Figure A1 compares the estimated CAMCEs for country with Germany set as the reference category, calculated across all individual covariate vectors in the sample. Our method discovers a considerable degree of heterogeneity in the CAMCEs whereas cjbart shows limited treatment effect variation for most countries. Under our model, the estimated heterogeneous effects are more strongly associated with predictors than cjbart. For example, our method finds a clear association, on average, between the estimated CAMCEs and prejudice or party identification, whereas cjbart does not.
Figure A1: Comparison of discovered heterogeneous effects (the conditional average marginal component effects or CAMCEs) between the proposed method and the BART-based method cjbart. In both plots the y-axis corresponds to values estimated, either by our method (right) or by cjbart (Robinson and Duch, 2024) (left). The plots show the estimated effect of Iraq as compared to the baseline of Germany. In the top figure, the x-axis and color corresponds to the categories of individuals based on the quartile of their Hispanic prejudice score. In the bottom figure, the x-axis and color corresponds to party ID.
Figure A2 shows the distribution of CAMCEs for all countries. To simplify the visualization, we subset party ID to strong Republicans, strong Democrats, and Independent/other.
Figure A2: In both plots, the y-axis corresponds to the estimated values, either based on our method (right) or based on the method of Robinson and Duch (2024) (left), for the effect of a given country relative to the baseline of Germany. In the top figure, we color code based on quartile for the Hispanic prejudice score. In the bottom figure, we reduce the sample to those who identify as “Strong Republican”, “Strong Democrat”, or “Independent/Other” and color code by party ID.
Supplementary Material C Proof of Proposition 1
To prove Proposition 1, we provide a more general result for one-parameter exponential family distributions, which include the specific Bernoulli result in the main text as a special case. We consider a random variable $Y$ that is assumed to follow a single-parameter exponential family distribution with canonical parameter $\theta$ and $\mu=d\psi(\theta)/d\theta=\psi^{\prime}(\theta)$. Since $\mu$ is monotone in $\theta$, we index the density $f$ using $\mu$:
$f_{\mu}(y)=c(y)\exp\left(y\theta-\psi(\theta)\right).$
The maximum likelihood estimate of the mean $\hat{\mu}$ given $N$ observations $\{y_{i}\}_{i=1}^{N}$ from $Y$ is the sample average, $\frac{1}{N}\sum_{i=1}^{N}y_{i}=\hat{\mu}$, and the corresponding estimate of the canonical parameter is $\hat{\theta}$.
Proposition C.1 states that maximally heterogeneous groups in terms of Kullback-Leibler (KL) divergence of potential outcomes is equivalent to maximizing the log-likelihood over groups and their centroids for any choice of single parameter exponential family $f$.
Proposition C.1.
Assume a partition of $N$ observations, indexed by $i\in\{1,\cdots,N\}$, into $K$ groups whose memberships $Z_{i}\in\{1,\cdots,K\}$ are denoted by $\mathcal{Z}$. Define the estimated within-group average outcome under treatment $\bm{t}$ for group $k$ and the estimated overall average outcome as $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})=\sum_{i=1}^{N}I\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}Y_{i}/\sum_{i=1}^{N}I\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}$ and $\widehat{\overline{Y}}(\bm{t})=\sum_{i=1}^{N}I\{\bm{T}_{i}=\bm{t}\}Y_{i}/\sum_{i=1}^{N}I\{\bm{T}_{i}=\bm{t}\}$, respectively.
Then, maximally heterogeneous groups in the terms of the Kullback-Leibler (KL) divergence of potential outcomes can be found by maximizing the log-likelihood function over the group membership and the centroids of groups, i.e.,
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)\right\}=\operatornamewithlimits{argmax}_{\mathcal{Z}}\sum_{k=1}^{K}\sup_{\zeta_{k}}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\log f_{\zeta_{k}}(Y_{i})$
where $\mathrm{KL}(\mu_{1},\mu_{2})$ indicates the KL divergence between two single-parameter exponential family distributions with means $\mu_{1}$ and $\mu_{2}$ is defined as (Hastie, 1987):
$\mathrm{KL}(\mu_{1},\mu_{2})=\mathbb{E}_{f_{\mu_{1}}(Y)}\left[\log f_{\mu_{1}}(Y)-\log f_{\mu_{2}}(Y)\right]=\left(\theta_{1}-\theta_{2}\right)\mu_{1}-\left[\psi(\theta_{1})-\psi(\theta_{2})\right].$
To prove this proposition, we use Lemma C.1 which decomposes the total deviance of the observed data into the between and within components as in $k$-means (Everitt et al., 2011, ch. 5). This generalizes the standard Gaussian result (see Chi, Chi and Baraniuk, 2016).
Lemma C.1 (Deviance Decomposition for Exponential Family).
Define the deviance for a single observation $y$ as follows:
$D(y,\mu)=2\left[\log f_{y}(y)-\log f_{\mu}(y)\right]$
and the total deviance of the observed data when evaluated at the maximum likelihood estimate for each treatment $\bm{t}$—the sample average $\widehat{\overline{Y}}(\bm{t})$ given randomization of $\bm{T}_{i}$—as follows
$D_{\mathrm{Total}}=\sum_{i=1}^{N}\sum_{\bm{t}\in\mathcal{T}}D\left(Y_{i}(\bm{t}),\widehat{\overline{Y}}(\bm{t})\right)\bm{1}\{\bm{T}_{i}=\bm{t}\}=\sum_{i=1}^{N}D\left(Y_{i},\widehat{\overline{Y}}(\bm{T}_{i})\right),$
where $\widehat{\overline{Y}}(\bm{t})=\sum_{i=1}^{N}\bm{1}\{\bm{T}_{i}=\bm{t}\}Y_{i}/\sum_{i=1}^{N}\bm{1}\{\bm{T}_{i}=\bm{t}\}$. Then, for any partition $\mathcal{Z}$ of the observations into $K$ groups, $D_{\mathrm{Total}}$ can be decomposed as follows:
$D_{\mathrm{Total}}=\underbrace{\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\cdot 2~{}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z}),\widehat{\overline{Y}}(\bm{T}_{i})\right)}_{=D_{\mathrm{Between}}}+\underbrace{\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}D\left(Y_{i}(\bm{T}_{i}),\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\right)}_{=D_{\mathrm{Within}}}$
where $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}Y_{i}/N_{k}(\bm{t};\mathcal{Z})$ and $N_{k}(\bm{t};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}$.
Proof.
Define $\hat{\bar{\theta}}(\bm{t})$, $\hat{\theta}_{k}(\bm{t};\mathcal{Z})$ and $\theta_{i}(\bm{t})$ as the canonical parameters associated with, respectively, means $\widehat{\overline{Y}}(\bm{t})$, $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})$, $Y_{i}$ where $\theta_{i}(\bm{t})$ is used to define a saturated model for $Y_{i}(\bm{t})$. The result is proved below by re-arranging $D_{\mathrm{Total}}$.
$D_{\mathrm{Total}}=$ $\sum_{i=1}^{N}\sum_{\bm{t}\in\mathcal{T}}\bm{1}\{\bm{T}_{i}=\bm{t}\}\cdot 2\left[\left(\theta_{i}(\bm{t})-\hat{\bar{\theta}}(\bm{t})\right)Y_{i}(\bm{t})-\left(\psi(\theta_{i}(\bm{t}))-\psi(\hat{\bar{\theta}}(\bm{t}))\right)\right]$
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{\bm{t}\in\mathcal{T}}\bm{1}\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}\cdot 2\left(\theta_{i}(\bm{t})-\hat{\bar{\theta}}(\bm{t})+\hat{\theta}_{k}(\bm{t};\mathcal{Z})-\hat{\theta}_{k}(\bm{t};\mathcal{Z})\right)Y_{i}(\bm{t})$
$-\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{\bm{t}\in\mathcal{T}}\bm{1}\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}\cdot 2\left(\psi(\theta_{i}(\bm{t}))-\psi(\hat{\bar{\theta}}(\bm{t}))+\psi(\hat{\theta}_{k}(\bm{t};\mathcal{Z}))-\psi(\hat{\theta}_{k}(\bm{t};\mathcal{Z}))\right)$
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{\bm{t}\in\mathcal{T}}\bm{1}\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}\cdot 2\hat{\zeta}_{k}(\bm{t};\mathcal{Z})\left[\left(\hat{\theta}_{k}(\bm{t};\mathcal{Z})-\hat{\bar{\theta}}(\bm{t})\right)-\left(\psi(\hat{\theta}_{k}(\bm{t};\mathcal{Z}))-\psi(\hat{\bar{\theta}}(\bm{t}))\right)\right]$
$+\sum_{k=1}^{K}\sum_{i=1}^{N}\sum_{\bm{t}\in\mathcal{T}}\bm{1}\{Z_{i}=k,\bm{T}_{i}=\bm{t}\}D\left(Y_{i}(\bm{t}),\hat{\zeta}_{k}(\bm{t};\mathcal{Z})\right)$
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\cdot 2~{}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z}),\widehat{\overline{Y}}(\bm{T}_{i})\right)+\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}D\left(Y_{i}(\bm{T}_{i}),\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\right)$
where the simplification of $D_{\mathrm{Between}}$ follows from noting that $\sum_{i=1}^{N}Y_{i}(\bm{t})\bm{1}\{\bm{T}_{i}=\bm{t},Z_{i}=k\}=N_{k}(\bm{t};\mathcal{Z})\hat{\zeta}_{k}(\bm{t})=\sum_{i=1}^{N}\bm{1}\{\bm{T}_{i}=\bm{t},Z_{i}=k\}\hat{\zeta}_{k}(\bm{t})$ by definition. ∎
Proof of Proposition C.1.
Given Lemma C.1, maximizing $D_{\mathrm{Between}}$ over $\mathcal{Z}$ is equivalent to minimizing $D_{\mathrm{Within}}$ over $\mathcal{Z}$. Then, $D_{\mathrm{Between}}$ can be divided by two to obtain the left-hand side of the proposition. The right-hand side of the proposition is derived as follows. Minimizing the deviance is equivalent to maximizing the log-likelihood, i.e.,
$\operatornamewithlimits{argmin}_{\mathcal{Z}}D_{\mathrm{Within}}=\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\log f_{\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})}(Y_{i}(\bm{T}_{i}))\right\}.$
This can be written as a two-level optimization problem, noting that $Y_{i}=Y_{i}(\bm{T}_{i})$ by the consistency assumption and that for fixed $\mathcal{Z}$, the maximum likelihood estimate of $\zeta_{k}(\bm{t})$ is $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})$, i.e., the within-group observed average.
$\operatornamewithlimits{argmin}_{\mathcal{Z}}D_{\mathrm{Within}}$ $=\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sup_{\zeta_{k}}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\log f_{\zeta_{k}(\bm{T}_{i})}(Y_{i})\right\}$
Finally, Proposition 1 in the main text uses the Bernoulli likelihood for $f$ and is shown below.
$\operatornamewithlimits{argmin}_{\mathcal{Z}}D_{\mathrm{Within}}=\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sup_{\{\zeta_{k}\}}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\left[Y_{i}\log\zeta_{k}(\bm{T}_{i})+\{1-Y_{i}\}\log\{1-\zeta_{k}(\bm{T}_{i})\}\right]\right\}$
Supplementary Material D Proof of Proposition 2
As before, we prove a more general result using the one-parameter exponential family distributions.
Proposition D.1 (Finding maximally heterogeneous groups with moderators).
Suppose we extend the setting of Proposition C.1 and additionally model the conditional probability of each individual’s group membership given categorical moderators $\{\pi_{k}(\bm{X}_{i})\}_{k=1}^{K}$. Then, maximally heterogeneous groups in terms of KL divergence of potential outcomes with the entropy of group membership probabilities as a penalty term can be found by maximizing the log-likelihood function of the extended model,
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)-\sum_{i=1}^{N}H(\{\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z})\}_{k=1}^{K})\right\}$
=\ = $\operatornamewithlimits{argmax}_{\mathcal{Z}}\sum_{k=1}^{K}\sup_{\zeta_{k},\pi_{k}}\sum_{i=1}^{N}\mathbf{1}\{Z_{i}=k\}\left[\log f_{\zeta_{k}}(Y_{i})+\log\pi_{k}(\bm{X}_{i})\right]$
where $H(\{p_{k}\}_{k=1}^{K})=-\sum_{k=1}^{K}p_{k}\log p_{k}$ (by convention, if $p_{k}=0$, then $p_{k}\log p_{k}=0$) is the entropy, and $\hat{\pi}_{k}(\bm{x};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{X}_{i}=\bm{x}\}/\sum_{i=1}^{N}\bm{1}\{\bm{X}_{i}=\bm{x}\}$ and $\hat{\zeta}_{k}(\bm{t};\mathcal{Z})$ are the maximizers of the log-likelihood function of the right hand side of the above equation given $\mathcal{Z}$.
To prove this proposition, we use Lemma D.1.
Lemma D.1 (Entropy of Groups with Respect to Moderators).
Define the set of observed categorical moderator values as $\mathcal{X}$ with $N(\bm{x})=\sum_{i=1}^{N}\bm{1}\{\bm{X}_{i}=\bm{x}\}$ and $N_{k}(\bm{x};\mathcal{Z})=\sum_{i=1}^{N}\bm{1}\{Z_{i}=k,\bm{X}_{i}=\bm{x}\}$. Given $\mathcal{Z}$, the entropy of group membership probabilities given moderators, weighted by the frequency of the moderators, is defined as follows:
$H(\mathcal{Z})=\sum_{\bm{x}\in\mathcal{X}}N(\bm{x})H(\{\hat{\pi}_{k}(\bm{x};\mathcal{Z})\}_{k=1}^{K}).$
Then, $H(\mathcal{Z})$ can be expressed in the following two equivalent ways:
$H(\mathcal{Z})=\sum_{i=1}^{N}H(\{\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z})\}_{k=1}^{K})=-\sum_{i=1}^{N}\sum_{k=1}^{K}\bm{1}\{Z_{i}=k\}\log\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z}).$
Proof.
The first expression follows by noting that the summation merely counts the number of times each $\bm{x}$ appears. The second expression is derived below by re-arranging $H(\mathcal{Z})$,
$-H(\mathcal{Z})$ $=\sum_{\bm{x}\in\mathcal{X}}\sum_{k=1}^{K}N(\bm{x})\hat{\pi}_{k}(\bm{x};\mathcal{Z})\log\hat{\pi}_{k}(\bm{x};\mathcal{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\bm{1}\{Z_{i}=k\}\log\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z}),$
where the last equality follows because $N(\bm{x})\hat{\pi}_{k}(\bm{x};\mathcal{Z})=N_{k}(\bm{x};\mathcal{Z})$ and it counts the number of times each combination of $(k,\bm{x})$ appears. ∎
Next, to prove Proposition D.1, we note that for any $\mathcal{Z}$, the KL divergence is equal to the log-likelihood evaluated at the maximum likelihood estimates plus a constant that does not depend on $\mathcal{Z}$ (see the definition of $D_{\mathrm{Total}}$):
$\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\mathrm{KL}\left(\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})\|\widehat{\overline{Y}}(\bm{T}_{i})\right)\ =\ \sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\log_{\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})}(Y_{i}(\bm{T}_{i}))+\mathrm{const.}$
Adding the negative of group-moderator entropy $H(\mathcal{Z})$ to both sides and taking the maximum over $\mathcal{Z}$ gives the left-hand side of Proposition D.1. The equivalent right-hand side, using Lemma D.1 can be expressed as:
$\operatornamewithlimits{argmax}_{\mathcal{Z}}\left\{\sum_{k=1}^{K}\sum_{i=1}^{N}\bm{1}\{Z_{i}=k\}\left[\log f_{\hat{\zeta}_{k}(\bm{T}_{i};\mathcal{Z})}(Y_{i}(\bm{T}_{i}))+\log\hat{\pi}_{k}(\bm{X}_{i};\mathcal{Z})\right]\right\}.$
As in the proof of Proposition C.1, observing that $Y_{i}=Y_{i}(\bm{T}_{i})$ by the consistency assumption and writing the above equation as two-level optimization problem over $\zeta_{k}$ and $\pi_{k}$ establishes Proposition D.1. This follows by noting that for a fixed $\mathcal{Z}$, the maximum likelihood estimate of $\pi_{k}(\bm{x})$ is $\hat{\pi}_{k}(\bm{x})$ and the estimate of $\zeta_{k}(\bm{t})$ as $\hat{\zeta}_{k}(\bm{t})$ is unchanged as the optimization problem is separable. In addition, using the Bernoulli likelihood for $f$ gives Proposition 2 in the main text. ∎
Supplementary Material E Inclusion of Higher Order Interactions
Here we illustrate how the model and regularization penalties in Section 3.3 can be extended to include higher order interactions in a straightforward manner. We show below the model including all higher order interactions, and including only a subset is direct.
Let $\mathcal{J}=\{1,\dots,J\}$ be the set of $J$ factors and let $\mathcal{T}$ be the set of all possible assignments on the $\mathcal{J}$ factors. Then our model for $\psi_{k}(\bm{T}_{i})$ with all interactions among factors is
$\psi_{k}(\bm{T}_{i})$ $\ =\ \mu+\sum_{j=1}^{J}\sum_{l=0}^{L_{j}-1}\mathbf{1}\{T_{ij}=l\}\beta^{j}_{kl}+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\sum_{l=0}^{L_{j}-1}\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}\beta^{jj^{\prime}}_{kll^{\prime}}$
$\qquad\ +\ \cdots+\sum_{\bm{t}\in\mathcal{T}}\mathbf{1}\{\bm{T}_{i}=\bm{t}\}\beta^{12\cdots K}_{k\bm{t}}$
$\ =\ \mu+\tilde{\bm{T}}_{i}^{\top}\bm{\beta}_{k}.$
In the above formulation, $\beta^{12\cdots K}_{k\bm{t}}$ is the $K$-way interaction coefficient in cluster $k$ for assignment $\bm{t}$.
Let $\mathcal{T}_{-j}$ be the set of all possible assignments on the $\mathcal{J}$ factors except for factor $j$. With some slight notation abuse by letting $\beta^{12\cdots K}_{kl\bm{t}_{-j}}$ be the $K$-way interaction coefficient in cluster $k$ for assignment $l$ for $j$ and $\bm{t}_{j}$ for the other $J-1$ factors, the ANOVA-type sum-to-zero constraints extend as follows:
$\sum_{l=0}^{L_{j}-1}\beta^{j}_{kl}=0,\sum_{l=0}^{L_{j}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\beta^{jj^{\prime}}_{kll^{\prime}}=0,\dots,\sum_{l=0}^{L_{j}-1}\beta^{12\cdots K}_{kl\bm{t}_{-j}}=0$ (A1)
for $j,j^{\prime}=1,2,\ldots,J$ with $j^{\prime}>j$ and for all $\bm{t}_{-j}\in\mathcal{T}_{-j}$. We write them compactly as,
[…]
where each row of $\bm{C}^{\top}\bm{\beta}_{k}$ corresponds to one of the constraints given in Equation (A1).
For the structured sparsity, we have penalties of the form
$\sum_{j=1}^{J}\sum_{l_{j}=1}^{L_{j}}\sum_{l_{j}^{\prime}>l_{j}}^{L_{j}}\sqrt{(\beta^{j}_{l_{j}}-\beta^{j}_{l_{j}^{\prime}})^{2}+\sum_{j^{\prime}\neq j}\sum_{l_{j^{\prime}}=1}^{L_{j^{\prime}}}(\beta^{jj^{\prime}}_{l_{j}l_{j^{\prime}}}-\beta^{j{j^{\prime}}}_{l_{j}^{\prime}l_{j^{\prime}}})^{2}+\cdots+\sum_{\bm{t}_{-j}\in\mathcal{T}_{-j}}(\beta^{12\cdots K}_{l_{j}\bm{t}_{-j}}-\beta^{12\cdots K}_{l_{j}^{\prime}\bm{t}_{-j}})^{2}}$
This will have $\sum_{j=1}^{J}L_{j}(L_{j}-1)/2$ terms, $L_{j}(L_{j}-1)/2$ terms for the $j$th factor.
For illustration, consider a simple example with one group and three factors—factor one has three levels, factor two has two levels, and factor three has two levels. In this case, our penalty contains 5 terms,
$\begin{split}&\sum_{l_{1}=1}^{L_{1}}\sum_{l_{1}^{\prime}>l_{1}}^{L_{1}}\sqrt{(\beta^{1}_{l_{1}}-\beta^{1}_{l_{1}^{\prime}})^{2}+\sum_{l_{2}=1}^{L_{2}}(\beta^{12}_{l_{1}l_{2}}-\beta^{12}_{l_{1}^{\prime}l_{2}})^{2}+\sum_{l_{3}=1}^{L_{3}}(\beta^{13}_{l_{1}l_{3}}-\beta^{13}_{l_{1}^{\prime}l_{3}})^{2}+\sum_{l_{2}=1}^{L_{2}}\sum_{l_{3}=1}^{L_{3}}(\beta^{123}_{l_{1}l_{2}l_{3}}-\beta^{123}_{l_{1}^{\prime}l_{2}l_{3}})^{2}}\\ ~{}+~{}&\sum_{l_{2}=1}^{L_{2}}\sum_{l_{2}^{\prime}>l_{2}}^{L_{2}}\sqrt{(\beta^{1}_{l_{2}}-\beta^{2}_{l_{2}^{\prime}})^{2}+\sum_{l_{2}=1}^{L_{2}}(\beta^{12}_{l_{1}l_{2}}-\beta^{12}_{l_{1}l_{2}^{\prime}})^{2}+\sum_{l_{3}=1}^{L_{3}}(\beta^{23}_{l_{2}l_{3}}-\beta^{23}_{l_{2}^{\prime}l_{3}})^{2}+\sum_{l_{1}=1}^{L_{1}}\sum_{l_{3}=1}^{L_{3}}(\beta^{123}_{l_{1}l_{2}l_{3}}-\beta^{123}_{l_{1}l_{2}^{\prime}l_{3}})^{2}}\\ \\ ~{}+~{}&\sum_{l_{3}=1}^{L_{3}}\sum_{l_{3}^{\prime}>l_{3}}^{L_{3}}\sqrt{(\beta^{1}_{l_{3}}-\beta^{3}_{l_{3}^{\prime}})^{2}+\sum_{l_{3}=1}^{L_{2}}(\beta^{13}_{l_{1}l_{3}}-\beta^{12}_{l_{1}l_{3}^{\prime}})^{2}+\sum_{l_{1}=1}^{L_{3}}(\beta^{13}_{l_{1}l_{3}}-\beta^{23}_{l_{1}l_{3}^{\prime}})^{2}+\sum_{l_{1}=1}^{L_{1}}\sum_{l_{2}=1}^{L_{2}}(\beta^{123}_{l_{1}l_{2}l_{3}}-\beta^{123}_{l_{1}l_{2}l_{3}^{\prime}})^{2}}\\ \end{split}$
The first three terms encourages the pairwise fusion of the levels of factor one whereas the fourth encourages the fusion of the two levels of factor two and the fifth encourages the fusion of the two levels of factor three.
Using the sum of Euclidean norms of quadratic forms, we can write the penalty as
$||\bm{\beta}^{\top}\bm{F}_{1}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{2}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{3}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{4}\bm{\beta}||_{2}+\ ||\bm{\beta}^{\top}\bm{F}_{5}\bm{\beta}||_{2},$
[…]
More generally, for a fully interacted model we will have $\sum_{j=1}^{J}L_{j}(L_{j}-1)/2=G$ terms,
$\sum_{g=1}^{G}||\bm{\beta}^{\top}\bm{F}_{g}\bm{\beta}||_{2}.$
Supplementary Material F Propriety of the Structured Sparse Prior
The proof of propriety for the structured sparse prior used in our paper is an application of Theorem 1 established in Goplerud (2021) and is reproduced here.
Theorem 1 (Goplerud (2021)).
Consider the following structured sparse prior on $\bm{\beta}\in\mathbb{R}^{p}$ with regularization strength $\lambda>0$ penalizes $K$ linear constraints $\bm{d}_{k}$ and $L$ quadratic constraints $\bm{F}_{\ell}$ on the parameters where $\bm{F}_{\ell}$ is symmetric and positive semi-definite. The kernel of the prior is shown below.
$p(\bm{\beta})\propto\exp\left(-\lambda\left[\sum_{k=1}^{K}|\bm{d}_{k}^{\top}\bm{\beta}|+\sum_{\ell=1}^{L}\sqrt{\bm{\beta}^{\top}\bm{F}_{\ell}\bm{\beta}}\right]\right)$
Further define $\bm{D}^{\top}=[d_{1},\cdots,d_{K}]^{\top}$ and $\bar{\bm{D}}^{\top}=[\bm{D}^{\top},\bm{F}_{1},\cdots,\bm{F}_{L}]$. Then, for $\lambda>0$, the prior above is proper if and only if $\bar{\bm{D}}$ is full column rank.
In our specific case, we note that $K=0$, $L=G$, and $\lambda=\lambda\bar{\pi}_{k}^{\gamma}$. Prior propriety of $p(\bm{\beta}_{k}\mid\{\bm{\phi}_{k}\}_{k=2}^{K},\lambda)$, therefore, can be determined by empirically investigating whether $\bar{\bm{D}}$, i.e. the vertically stacked $\bm{F}_{\ell}$, is full column rank.
It is also possible to analytically show the propriety of the prior distribution in all cases considered in this paper. We focus on the case of $K=1$ and arbitrary $\lambda>0$ as the result follows automatically for the case in our paper.
Result F.1.
Assume a structured sparse prior for a factorial or conjoint design with $J$ factors each with $L_{j}$ levels where all pairwise interactions are included and levels of each factor are encouraged to be fused together (i.e. the model in the main text). The kernel of the prior is shown below where $\bm{F}_{g}$ are as defined in the main text.
$k(\bm{\beta})=\exp\left(-\lambda\sum_{g=1}^{G}\sqrt{\bm{\beta}^{\top}\bm{F}_{g}\bm{\beta}}\right)$
Assume that the linear sum-to-zero constraints $\bm{C}^{\top}\bm{\beta}=\bm{0}$ hold. Then, the structured sparse prior on the unconstrained $\tilde{\bm{\beta}}$ such that $\tilde{\bm{\beta}}\in\mathcal{N}(\bm{C}^{\top})$ is proper. Or, equivalently, the following result holds.
[…]
Proof.
Let $\mathcal{B}_{\bm{C}^{\top}}$ represent a basis for the linear constraints $\bm{C}^{\top}$. The integral for evaluating propriety can be written as,
[…]
Note that $\bm{F}_{g}$ can be expressed as a sum of $N_{g}$ outer products of $|\bm{\beta}|$-length vectors of the form $\bm{l}_{i}\in\{-1,0,1\}$ where $-1$ and $1$ correspond to the two terms that are fused together and all other elements are 00, i.e., $\bm{F}_{g}=\sum_{g^{\prime}=1}^{N_{g}}\bm{l}_{g^{\prime}}\bm{l}_{g^{\prime}}^{\top}$. Thus, one can define a matrix $\bm{Q}_{g}^{\top}=\left[\bm{l}_{1},\cdots,\bm{l}_{N_{g}}\right]$ such that $\bm{Q}_{g}^{\top}\bm{Q}_{g}=\bm{F}_{g}$, which allows us to rewrite $\tilde{k}(\tilde{\bm{\beta}})$ as:
$\tilde{k}(\tilde{\bm{\beta}})=\exp\left(-\lambda\sum_{g=1}^{G}\sqrt{\tilde{\bm{\beta}}^{\top}\left[\mathcal{B}_{\bm{C}^{\top}}\right]^{\top}\bm{Q}_{g}^{\top}\bm{Q}_{g}\mathcal{B}_{\bm{C}^{\top}}\tilde{\bm{\beta}}}\right).$
By applying Theorem 1 and noting that the null spaces of $\bm{A}^{T}\bm{A}$ and $\bm{A}$ are identical, the integral of $\tilde{k}(\tilde{\bm{\beta}})$ is finite if and only if $\bm{Q}\mathcal{B}_{\bm{C}^{\top}}$ is full column rank, where $\bm{Q}^{\top}=[\bm{Q}^{\top}_{1},\cdots,\bm{Q}^{\top}_{G}]$. We demonstrate this fact in two steps. First, there exists a permutation matrix $\bm{P}_{Q}$ such that $\bm{P}_{Q}\bm{Q}$ has a block diagonal structure with $J+1$ diagonal blocks. The first $J$ blocks corresponding to the main terms for each factor $j$ and the last block corresponds to all interaction terms. The null space of each block is spanned by the vector $\bm{1}$ as the corresponding block of $\bm{P}_{Q}\bm{Q}$ is a (transposed) orientated incidence matrix of a fully connected graph. Thus, the null space of $\bm{P}_{Q}\bm{Q}$, and hence $\bm{Q}$, is spanned by the $J+1$ columns of a block diagonal matrix with $\bm{1}$ on each block. Second, consider the linear constraints $\bm{C}^{\top}\bm{\beta}=\bm{0}$. The only vector to satisfy this constraint and lie in the null space of $\bm{Q}$ must be $\bm{0}$ as, for each block, the only vector proportional to $\bm{1}$ and satisfying the corresponding sum-to-zero constraints must be $\bm{0}$. Thus, $\bm{Q}\mathcal{B}_{\bm{C}^{\top}}$ is full column rank and the prior is proper. ∎
Supplementary Material G Derivations for the Basic Model
This section derives a number of results for the basic model. It first restates the main results concerning the elimination of the linear constraints $\bm{C}^{\top}\bm{\beta}_{k}=\bm{0}$. Then, it derives the Expectation Maximization algorithm, our measure of degrees of freedom, and some additional computational improvements used to accelerate estimation. In the following, we use $\tilde{\bm{T}}_{i}$ to denote the corresponding vector of indicators for whether certain treatments or interactions are present (i.e. stacking all $\bm{1}\{T_{ij}=l\}$, etc. from Equation A3). In addition, we use $\psi_{ik}$ to indicate the linear predictor for observation $i$ and group $k$.
G.1 Removing the Linear Constraints
The inference problem in the main text is presented as an optimization problem subject to linear constraints on the coefficients $\bm{\beta}_{k}$. Inference is noticeably easier if these are eliminated via a transformation of the problem to a lower-dimensional one by noting that $\bm{\beta}_{k}$ must lie in the null space of the constraint matrix $\bm{C}^{\top}$ (see, e.g., Lawson and Hanson 1974, ch. 20). Define $\tilde{\bm{\beta}}_{k}=\left(\mathcal{B}_{\bm{C}^{\top}}^{\top}\mathcal{B}_{\bm{C}^{\top}}\right)^{-1}\mathcal{B}_{\bm{C}^{\top}}^{\top}\bm{\beta}_{k}$ where $\mathcal{B}_{\bm{C}^{\top}}$ is a basis for the null space of $\bm{C}^{\top}$. The problem can thus be solved in terms of the unconstrained $\tilde{\bm{\beta}}_{k}\in\mathbb{R}^{p-\mathrm{rank}(\bm{C}^{\top})}$ given appropriate adjustment of the treatment design vectors, $\tilde{\tilde{\bm{T}}}_{i}=\mathcal{B}_{\bm{C}^{\top}}\tilde{\bm{T}}_{i}$, penalty matrices, $\tilde{\bm{F}}_{g}=\mathcal{B}_{\bm{C}^{\top}}^{\top}\bm{F}_{g}\mathcal{B}_{\bm{C}^{\top}}$, and linear predictor, $\psi_{i,k}=\left[\tilde{\tilde{\bm{T}}}_{i}\right]^{\top}\tilde{\bm{\beta}}_{Z_{i}}+\mu$. Once the algorithm convergences, the constrained parameters can be recovered by noting $\bm{\beta}_{k}=\mathcal{B}_{\bm{C}^{\top}}\tilde{\bm{\beta}}_{k}$.
Given the similarity of the unconstrained and constrained problems and for notational simplicity, we present all results herein dropping the second “tilde” notation on $\tilde{\bm{T}}_{i}$ and the “tilde” on $\bm{\beta}_{k}$ and note that, once estimated, $\tilde{\bm{\beta}}_{k}$ is projected back into the original space for the reported coefficients, average marginal component effects, etc. The results of Appendix J on approximating $\tilde{\bm{\beta}}_{k}$ as multivariate Gaussian imply that $\bm{\beta}_{k}$ will have a (singular) multivariate Gaussian distribution.
G.2 Expectation Maximization Algorithm
This section considers inference after removing the linear constraints as discussed in the prior subsection. Algorithm A1 summarizes our approach to maximizing Equation (3.10). Each iteration of our AECM algorithm involves two cycles where the data augmentation scheme enables iterative updating of the treatment effect parameters $\bm{\beta}$ and moderators $\bm{\phi}$. $\bm{\theta}$ collects both sets of parameters.
Algorithm A1 AECM Algorithm for Estimating $\bm{\theta}$
Set Hyper-Parameters: $K$ (groups), $\lambda$, $\sigma^{2}_{\phi}$, $\gamma$ (prior strength), $\epsilon_{1},\epsilon_{2}$ (convergence criteria), $T$ (number of iterations)
Initialize Parameters: $\bm{\theta}^{(0)}$, i.e. $\bm{\beta}^{(0)}$ and $\bm{\phi}^{(0)}$; Appendix G.5 provides details.
For iteration $t\in\{0,\cdots,T-1\}$
     Cycle 1: Update $\bm{\beta}$
     1a. $E$-Step: Find the conditional distributions of $\{Z_{i},\omega_{i}\}_{i=1}^{N}$ and $\{\{\tau^{2}_{gk}\}_{g=1}^{G}\}_{k=1}^{K}$ given $\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}$ and $\bm{\theta}^{(t)}$ (Eq. (A1)). Derive $Q_{\beta}(\bm{\beta},\bm{\theta}^{(t)})$ (Eq. (G.2.1)).
     1b. $M$-Step: Set $\bm{\beta}^{(t+1)}$ such that $Q_{\beta}(\bm{\beta}^{(t+1)},\bm{\theta}^{(t)})\geq Q_{\beta}(\bm{\beta}^{(t)},\bm{\theta}^{(t)})$
     Cycle 2: Update $\bm{\phi}$
     2a. $E$-Step: Find $p(Z_{i}=k\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\beta}^{(t+1)},\bm{\phi}^{(t)})$. Derive $Q_{\phi}(\bm{\phi},\{\bm{\beta}^{(t+1)},\bm{\phi}^{(t)}\})$ (Eq. (G.2.2)).
     2b. $M$-Step: Set $\bm{\phi}^{(t+1)}$ such that $Q_{\phi}(\bm{\phi}^{(t+1)},\{\bm{\beta}^{(t+1)},\bm{\phi}^{(t)}\})\geq Q_{\phi}(\bm{\phi}^{(t)},\{\bm{\beta}^{(t+1)},\bm{\phi}^{(t)}\})$
     Check Convergence
     3. Stop if $\log p\left(\bm{\theta}^{(t+1)}|\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)-\log p\left(\bm{\theta}^{(t)}|\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N}\right)<\epsilon_{1}$ (Eq. (3.10)) or
$||\bm{\theta}^{(t+1)}-\bm{\theta}^{(t)}||_{\infty}<\epsilon_{2}$.
G.2.1 Updating Treatment Effect Parameters
We begin with the cycle of the AECM algorithm for updating $\{\bm{\beta}_{k}\}_{k=1}^{K}$ and $\mu$ given $\{\bm{\phi}_{k=2}^{K}\}$. To update $\bm{\beta},\mu$, our data augmentation strategy requires three types of missing data. First, we use the standard group memberships of each unit $i$ for inference in finite mixtures, i.e., $Z_{i}\in\{1,\cdots,K\}$. We also include two other types of data augmentation that result in a closed-form update. We use Polya-Gamma augmentation ($\omega_{i}$; Polson, Scott and Windle 2013) for the logistic likelihood and data augmentation on the sparsity-inducing penalty ($\tau^{2}_{gk}$; see, e.g., Figueiredo 2003; Polson and Scott 2011; Ratkovic and Tingley 2017; Goplerud 2021) yielding
$p(Y_{i},\omega_{i}\mid Z_{i},\bm{X}_{i},\bm{T}_{i})\ \propto\ \frac{1}{2}\exp\left\{\left(Y_{i}-\frac{1}{2}\right)\psi_{Z_{i}}(\bm{T}_{i})-\frac{\omega_{i}}{2}\left[\psi_{Z_{i}}(\bm{T}_{i})\right]^{2}\right\}f_{PG}(\omega_{i}\mid 1,0),$ (A1a)
$p(\bm{\beta}_{k},\{\tau^{2}_{gk}\}_{g=1}^{G}\mid\lambda,\{\bm{\phi}_{k}\})\ \propto\ \exp\left\{-\frac{1}{2}\bm{\beta}_{k}^{\top}\left(\sum_{g=1}^{G}\frac{\bm{F}_{g}}{\tau^{2}_{gk}}\right)\bm{\beta}_{k}\right\}\prod_{g=1}^{G}\tau_{gk}^{-1}\exp\left\{-\frac{\left(\lambda\bar{\pi}_{k}\right)^{2}}{2}\cdot\tau^{2}_{gk}\right\},$ (A1b)
where $f_{PG}(\cdot\mid b,c)$ represents the Polya-Gamma distribution with parameters $(b,c)$ and $Z_{i}\sim\mathrm{Multinomial}\left(1,\bm{\pi}_{i}\right)$ with the $k$th element of $\bm{\pi}$ equal to $\pi_{k}(\bm{X}_{i})$. Note that $\bm{\beta}$ only enters Equation (A1) via a quadratic form. The first cycle of the AECM algorithm involves, therefore, maximizing the following function with respect to $\bm{\beta}$ given $\bm{\theta}^{(t)}$.
$Q_{\beta}\left(\bm{\beta},\bm{\theta}^{(t)}\right)\ =$ $\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{E}[\mathbf{1}\{Z_{i}=k\}]\left\{\left(Y_{i}-\frac{1}{2}\right)\psi_{k}(\bm{T}_{i})-\mathbb{E}[\omega_{i}\mid Z_{i}=k]\frac{\left[\psi_{k}(\bm{T}_{i})\right]^{2}}{2}\right\}$
$+\sum_{k=1}^{K}-\frac{1}{2}\bm{\beta}_{k}^{\top}\left[\sum_{g=1}^{K}\bm{F}_{g}\cdot\mathbb{E}[1/\tau^{2}_{gk}]\right]\bm{\beta}_{k}+\text{const.}$. (A2)
where all expectations are taken over the conditional distribution of the missing data given the current parameter estimates. We note that the $E$-Step involves computing $p(\{\omega_{i},Z_{i}\},\{1/\tau^{2}_{gk}\}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\},\bm{\theta}^{(t)})$ which factorizes into, respectively, a collection of Polya-Gamma (PG), categorical, and Inverse-Gaussian random variables. Their conditional distributions are shown below,
$p(\tau^{-2}_{gk}\mid\bm{\theta})$ $\sim\mathrm{InverseGaussian}\left(\frac{\lambda}{\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}},\quad\lambda^{2}\right),$ (A3a)
$p(Z_{i}=k\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\theta})$ $\propto p_{ik}^{Y_{i}}(1-p_{ik})^{1-Y_{i}}\pi_{ik};\quad p_{ik}=\frac{\exp(\psi_{ik})}{1+\exp(\psi_{ik})},$ (A3b)
$p(\omega_{i}\mid Z_{i}=k,\bm{X}_{i},\bm{T}_{i},\bm{\theta})$ $\sim\mathrm{PG}\left(1,\psi_{ik}\right),$ (A3c)
as well as the relevant expectations needed in $Q_{\beta}(\bm{\beta},\bm{\theta})$,
$\mathbb{E}\left[\tau^{-2}_{gk}\right]=\frac{\lambda}{\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}},$ (A4a)
$\mathbb{E}[z_{ik}]=\mathbb{E}\left[\bm{1}\{Z_{i}=k\}\right]=\frac{p_{ik}^{Y_{i}}(1-p_{ik})^{1-Y_{i}}\pi_{ik}}{\sum_{\ell=1}^{K}p_{i\ell}^{Y_{i}}(1-p_{i\ell})^{1-Y_{i}}\pi_{i\ell}},$ (A4b)
$\mathbb{E}[\omega_{i}\mid Z_{i}=k]=\frac{1}{2\psi_{ik}}\tanh\left(\frac{\psi_{ik}}{2}\right).$ (A4c)
Note that as $\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}$ approaches zero, $\mathbb{E}[\tau^{-2}_{gk}]$ approaches infinity. To prevent numerical instability, we rely on the strategy in Goplerud (2021) (inspired by Polson and Scott 2011) where once it is sufficiently small, e.g. below $10^{-4}$, and thus the restriction is almost binding, we ensure that restriction holds in all future iterations. We do so by adding a quadratic constraint $\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}=0$. This implies that $\bm{\beta}_{k}$ lies in the null space of $\bm{F}_{g}$ and thus with an additional transformation, it can be removed and the problem be solved in an unconstrained space with a modified design.
[…]
One could reply on a generalized EM algorithm where $Q_{\beta}$ is improved versus maximized for computational reasons, e.g. by using a conjugate gradient solver initialized at $\check{\bm{\beta}}^{(t)}$.
G.2.2 Updating Moderator Parameters
To update the moderator parameters $\bm{\phi}$, we use the second cycle of the AECM algorithm where only the $Z_{i}$ are treated as missing data. The $E$-step involves recomputing the group membership probabilities, i.e., $p(Z_{i}\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\beta}^{(t+1)},\bm{\phi}^{(t)})$, given the updates in the first cycle. The implied $Q$-function is shown below,
$Q_{\phi}(\bm{\phi},\{\bm{\beta}^{(t+1)},\bm{\phi}^{(t)}\})\ =$ $\sum_{k=1}^{K}\left[\sum_{i=1}^{N}\mathbb{E}[\mathbf{1}\{Z_{i}=k\}]\log\pi_{k}(\bm{X}_{i})\right]$
$+\sum_{k=1}^{K}\left[m\gamma\log\bar{\pi}_{k}-\lambda\bar{\pi}_{k}^{\gamma}\sum_{g=1}^{G}\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}\right]+\log p(\{\bm{\phi}_{k}\}_{k=2}^{K}),$ (A5)
where $\pi_{k}(\bm{X}_{i})$ and $\bar{\pi}_{k}=\sum_{i=1}^{N}\pi_{k}(\bm{X}_{i})/N$ are functions of $\bm{\phi}_{k}$. Note that if $\gamma=0$, this simplifies to a multinomial logistic regression with $\{\mathbb{E}[\bm{1}\{Z_{i}=k\}]\}_{k=1}^{K}$ as the outcome. We perform the $M$-Step using a standard optimizer (e.g., L-BFGS) to optimize $Q_{\phi}$ and thus obtain $\bm{\phi}^{(t+1)}$.
G.3 Classification Maximum Likelihood
If classification maximum likelihood approach is desired, despite statistical concerns about this procedure’s asymptotic bias (e.g., Bryant and Williamson 1978), it can be easily implemented by adapting the preceeding EM algorithm. Celeux and Govaert (1992) propose the “classification EM” algorithm in the spirit of how $k$-means classification is commonly implemented.
The adjustment proceeds as follows (Celeux and Govaert, 1992, p. 319): after conducting an $E$-step and obtaining $\tilde{\pi}_{k}(\bm{X}_{i},Y_{i},\bm{T}_{i};\bm{\theta})=\tilde{\pi}_{ik}=p(Z_{i}=k\mid Y_{i},\bm{X}_{i},\bm{T}_{i},\bm{\theta})$ for use in evaluating $Q_{\beta}$ and $Q_{\phi}$, perform a classification or “hard assignment”. That is, find $k^{*}_{i}=\operatornamewithlimits{argmax}_{k}\tilde{\pi}_{ik}$, i.e., the most probable cluster for observation $i$ given its observed data and $\bm{\theta}$. In the subsequent $M$-step, use a modified weight $c_{ik}=1$ if $k=k^{*}_{i}$ and otherwise $c_{ik}=0$ in lieu of $\tilde{\pi}_{ik}$.
G.4 Degrees of Freedom
Our procedure for estimating $\check{\bm{\beta}}^{(t)}$ appears similar to the results in Oelker and Tutz (2017) where complex regularization and non-linear models can be recast as a (weighted) ridge regression. Using that logic, we take the trace of the “hat matrix” implied by our algorithm at stationarity to estimate our degrees of freedom. We also adjust upwards the degrees of freedom by the number of moderator coefficients (e.g., Khalili 2010; Chamroukhi and Huynh 2019).
Equation (A6) shows our procedure where $\bm{\mathcal{R}}$ and $\check{\bm{\Omega}}$ contain expectations calculated at convergence. $p_{x}$ denotes the number of moderators, i.e. the dimensionality of $\bm{\phi}_{k}$. Before evaluating Equation (A6), for any two factor levels that are sufficiently close (e.g., $\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}<10^{-4}$), we assume they are fused together and consider it as an additional linear constraint on the parameter vector $\bm{\beta}_{k}$.
[…]
From this, we can calculate a BIC criterion. We seek to find the regularization parameter $\lambda$ that minimizes this criterion. To avoid the problems of a naive grid-search, we use Bayesian model-based optimization that attempts to minimize the number of function evaluations while searching for the value of $\lambda$ that minimizes the BIC (mlrMBO; Bischl et al. 2018). We find that with around fifteen model evaluations, the optimizer can usually find a near optimal value of $\lambda$.
G.5 Computational Improvements
While the algorithm above provides a valid way to locate a posterior mode, our estimation problem is complex and high-dimensional. Furthermore, given the complex posterior implied by mixture of experts models, we derived a number of computational strategies to improve convergence. We use the SQUAREM algorithm (Varadhan and Roland 2008). Our software provides the option to use a generalized EM algorithm to update $\bm{\beta}$ using a conjugate gradient approach and $\bm{\phi}$ using a few steps of L-BFGS.
We also outline a way to deterministically initialize the model to provide stability and, again, speed up estimation on large problems. To do this, we adapt the procedure from Murphy and Murphy (2020) for initializing mixture of experts: (i) initialize the groups using some (deterministic) procedure (e.g. spectral clustering on the moderators); (ii) using only the main effects, estimate an EM algorithm—possibly with hard assignment at the $E$-Step (CEM; Celeux and Govaert 1992); (iii) iterate until the memberships have stabilized. Use those memberships to initialize the model. This has the benefit of having a deterministic initialization procedure where the group membership is based on the moderators but guided by which grouping seem to have sensible treatment effects, at least for the main effects. Given the memberships, update $\bm{\beta}$ using a ridge regression and $\bm{\phi}$ using a ridge regression and take those values as $\bm{\beta}^{(0)}$ and $\bm{\phi}^{(0)}$.
Supplementary Material H Extensions to the Basic Model
As noted in the main text, there are five major extensions to the basic model that applied users might wish to include:
  1. 1.
    Repeated tasks (observations) for a single individual
  2. 2.
    A forced-choice conjoint experiment
  3. 3.
    Survey to weight the sample estimate to the broader population
  4. 4.
    Adaptive weights for each penalty
  5. 5.
    Latent overlapping groups
All can be easily incorporated into the proposed framework above. This section outlines the changes to the underlying model.
H.1 Repeated Observations
This modification notes that in factorial and conjoint experiments it is common for individuals to perform multiple tasks. Typically, the number of tasks $N_{i}$ is similar across individuals. The updated likelihood for a single observation $i$ is shown below; we show both the observed and complete case. $y_{im}$ represents the choice of person $i$ on task $m\in\{1,\cdots,N_{i}\}$; $p_{imk}$ is the probability of $Y_{im}=1$ if person $i$ was in group $k$, and $\tilde{\bm{T}}_{im}$ is the vector of treatment indicators for person $i$ on task $m$.
$L\left(\{Y_{im}\}_{m=1}^{N_{i}}\right)=\sum_{k=1}^{K}\pi_{ik}\left[\prod_{m=1}^{N_{i}}p_{imk}^{Y_{im}}(1-p_{imk})^{1-Y_{im}}\right];\quad p_{imk}=\frac{\exp(\psi_{imk})}{1+\exp(\psi_{imk})};\quad\psi_{imk}=\tilde{\bm{T}}_{im}^{\top}\bm{\beta}_{k}+\mu$ (A1)
$L^{c}(\{y_{im},\omega_{im}\}\mid Z_{i})=\prod_{t=1}^{N_{i}}\left[\frac{1}{2}\exp\left\{\left(Y_{im}-\frac{1}{2}\right)\psi_{i,Z_{i}}-\omega_{im}\frac{\psi_{im,Z_{i}}^{2}}{2}\right\}f_{PG}(\omega_{im}\mid 1,0)\right]$ (A2)
Note that because of the conditional independence of $(y_{it},\omega_{it})$ given $Z_{i}$ and the parameters, the major modifications to the EM algorithm is that the $E$-Step must account for all $t$ observations, i.e. the terms summed in Equation (A1). Some additional book-keeping is required in the code as the design of the treatments has $\sum_{i=1}^{N}N_{i}$ rows whereas the design of the moderators has $N$ rows. Repeated observations can be easily integrated into the uncertainty estimation procedure outlined below.
H.2 Forced Choice Conjoint Design
A popular design of a conjoint experiment is the forced choice design where the respondents are required to choose between two profiles. Therefore, the researcher does not observe an outcome for each profile separately, but rather a single outcome is observed for each pair indicating which is preferred. Egami and Imai (2019) show that this can be easily fit into the above framework with some adjustment. Specifically, the model is modified to difference the indicators of the treatment levels for the pair of profiles (subtracting, e.g., the levels of the profile presented on the left from those of the profile presented on the right). The intercept for this model can be interpreted as a preference for picking a profile presented in a particular location. With this modification, estimation proceeds as before.
H.3 Standardization Weights
An additional modification to the problem is to weight the penalty. This could be done for two reasons. First, there is an issue of the columns having different variances/Euclidean norms because of the different number of factor levels $L_{j}$. Second, it is popular to weight the penalty based on some consistent estimator (e.g. ridge regression) to improve performance and, in simpler models, can be shown to imply various oracle properties (e.g. Zou 2006). We leave the latter to future exploration.
Define $\xi_{gk}$ as a positive weight for the $g$-th penalty and the $k$-th group. The kernel of the penalty is modified to include them.
$\log p(\bm{\beta}_{k}\mid\lambda,\gamma,\{\bm{\phi}_{k}\})\propto-\lambda\bar{\pi}_{k}^{\gamma}\sum_{g=1}^{G}\xi_{gk}\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}$ (A3)
This has no implication on the rank of the stacked $\bm{F}_{g}$ (and thus the results in Appendix F) as they are all positive and thus only slightly modify the $E$-Step.
We employ weights in all of our analyses to account for the fact that different factors $j$ may have different number of levels $L_{j}$. We use a generalization of the weights in Bondell and Reich (2009) to the case of penalized differences. Specifically, consider the over-parameterized model in Appendix H.4 where the penalty can be written entirely on the differences $\bm{\delta}_{\mathrm{Main}}$, $\bm{\delta}_{\mathrm{Int}}$, $\bm{\delta}_{\mathrm{Main-Copy}}$. Note that each of those penalties has a simple (group) LASSO form and thus we adopt the approach in Lim and Hastie (2015) of weighting by the Frobenius norm of the associated columns in $\bm{T}_{\mathrm{LOG}}$, i.e. the over-parameterized design matrix. At slight abuse of notation, define $[\bm{T}_{\mathrm{LOG}}]_{g}$ as the columns of $\bm{T}_{\mathrm{LOG}}$ corresponding to the differences penalized in the (group) lasso $g$, the weight can be expressed as follows:
$\xi_{gk}=\frac{1}{\sqrt{N}}||~{}\left[\bm{T}_{\mathrm{LOG}}\right]_{g}||_{F}$
Ignoring the factor of $\sqrt{N}$, this exactly recovers the weight proposed in Bondell and Reich (2009) in the non-latent-overlapping non-interactive model of $(L_{j}+1)^{-1}\sqrt{N^{j}_{l}+N^{j}_{l^{\prime}}}$ where $N^{j}_{l}$, $N^{j}_{l^{\prime}}$ are the number of observations for factor $j$ in level $l$ and level $l^{\prime}$ that are being encouraged to fuse together by the penalty in group $g$.
H.4 Latent Overlapping Groups
One feature of the above approach is that our groups are highly overlapping. Yan and Bien (2017) suggest that, in this setting, a different formulation of the problem may result in superior performance (see also Lim and Hastie 2015). Existing work on the topic has focused on group LASSO penalties (e.g. $\bm{F}_{g}=\bm{I}$) and thus some modifications are needed for our purposes. To address this, we note that we can again recast our model in an equivalent fashion. Instead of penalizing $\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}$, we can penalize the vector of differences between levels as long as we also impose linear constraints to ensure that the original model is maintained.
Consider a simple example with two factors each with two levels $\{1,2\}$ and $\{A,B\}$. The relevant differences are defined such that $\delta^{j}_{1-2}=\beta^{j}_{1}-\beta^{j}_{2}$ and $\delta^{jj^{\prime}}_{(lm)-(l^{\prime}m^{\prime})}=\beta^{j}_{l,m}-\beta^{j^{\prime}}_{l^{\prime},m^{\prime}}$. The equivalent penalty can be imposed as follows:
$\begin{split}&\sqrt{\left(\delta^{j}_{1-2}\right)^{2}+\left(\delta^{jj^{\prime}}_{(1A)-(2A)}\right)^{2}+\left(\delta^{jj^{\prime}}_{(1B)-(2B)}\right)^{2}}=\sqrt{\bm{\delta}^{\top}\bm{\delta}};\quad\bm{\delta}=\left(\begin{array}[]{c}\delta^{j}_{1-2}\\ \delta^{jj^{\prime}}_{(1A)-(2A)}\\ \delta^{jj^{\prime}}_{(1B)-(2B)}\end{array}\right)\\ &\mathrm{such~{}that}~{}\left[\begin{array}[]{l}\delta^{j}_{1-2}\\ \delta^{jj^{\prime}}_{(1A)-(2A)}\\ \delta^{jj^{\prime}}_{(1B)-(2B)}\end{array}\right]=\left[\begin{array}[]{lll}\beta^{j}_{1}-\beta^{j}_{2}\\ \beta^{jj^{\prime}}_{1A}-\beta^{jj^{\prime}}_{2A}\\ \beta^{jj^{\prime}}_{1B}-\beta^{jj^{\prime}}_{2B}\end{array}\right]\end{split}$ (A4)
The latent overlapping group suggests a slight modification. In addition to the above penalization of the $\ell_{2}$ norm of the main and interactive differences, [^1] Note the related “hierarchical group LASSO” would add separate individual penalties for each of the interactions. It is easy to include that in our approach. it duplicates the main effect and penalizes it separately while ensuring that all effects maintain the accounting identities between the “latent” groups and the overall effect. Specifically, it modifies the above penalty to duplicate the column corresponding to $\delta^{j}_{1-2}$ and adds a new parameter $\delta^{j}_{(1-2)-\mathrm{Copy}}$.
$\begin{split}\sqrt{\bm{\delta}^{\top}\bm{\delta}}+|\delta^{j}_{(1-2)-\mathrm{Copy}}|\quad\mathrm{such~{}that}~{}\left[\begin{array}[]{l}\delta^{j}_{1-2}\\ \delta^{jj^{\prime}}_{(1A)-(2A)}\\ \delta^{jj^{\prime}}_{(1B)-(2B)}\end{array}\right]+\left[\begin{array}[]{l}\delta^{j}_{(1-2)-\mathrm{Copy}}\\ 0\\ 0\end{array}\right]=\left[\begin{array}[]{lll}\beta^{j}_{1}-\beta^{j}_{2}\\ \beta^{jj^{\prime}}_{1A}-\beta^{jj^{\prime}}_{2A}\\ \beta^{jj^{\prime}}_{1B}-\beta^{jj^{\prime}}_{2B}\end{array}\right]\end{split}$ (A5)
Scoping out to the full problem, define $\bm{\delta}_{\mathrm{Main}}$ as the main effect differences, e.g. $\delta^{j}_{1-2}$, and $\bm{\delta}_{\mathrm{Int}}$ as the interaction differences and $\bm{D}_{\mathrm{Main}}$ as the matrix such that $\bm{D}_{\mathrm{Main}}\bm{\beta}=\bm{\delta}_{\mathrm{Main}}$, and $\bm{D}_{\mathrm{Int}}$ as the corresponding matrix to create the vector of interactions. Define $\bm{\delta}_{\mathrm{Main}-g}$ as the sub-vector of $\bm{\delta}_{\mathrm{Main}-g}$ that corresponds to the (main) effect differences between levels $l$ and $l^{\prime}$ of factor $j$ penalized by $\bm{F}_{g}$ in the original notation. Similarly define $\bm{\delta}_{\mathrm{Int}-g}$ and $\bm{\delta}_{\mathrm{Main-Copy}-g}$.
$\begin{split}&p(\bm{\beta},\bm{\delta}_{\mathrm{Main}},\bm{\delta}_{\mathrm{Int}},\bm{\delta}_{\mathrm{Main-Copy}})=\sum_{g=1}^{G}\sqrt{\bm{\delta}_{\mathrm{Main}-g}^{T}\bm{\delta}_{\mathrm{Main}-g}+\bm{\delta}_{\mathrm{Int}-g}^{T}\bm{\delta}_{\mathrm{Int}-g}}+\sum_{g^{\prime}=1}^{G}\sqrt{\left[\bm{\delta}_{\mathrm{Main-Copy}-g}\right]^{2}}\\ &\mathrm{s.t.}\quad\left[\begin{array}[]{llll}\bm{C}^{\top}&\bm{0}&\bm{0}&\bm{0}\\ \bm{D}_{\mathrm{Main}}&-\bm{I}&\bm{0}&-\bm{I}\\ \bm{D}_{\mathrm{Int}}&\bm{0}&-\bm{I}&\bm{0}\end{array}\right]\left[\begin{array}[]{l}\bm{\beta}\\ \bm{\delta}_{\mathrm{Main}}\\ \bm{\delta}_{\mathrm{Int}}\\ \bm{\delta}_{\mathrm{Main-Copy}}\end{array}\right]=\bm{0}\end{split}$ (A6)
This also requires a modification of the design matrix $\tilde{\bm{T}}$ to ensure that (i) its dimensionality conforms with the expanded parameter vector and (ii) that for any choice of the expanded parameter that satisfies the constraints, the linear predictor for all observation (and thus the likelihood) is unchanged. Consider first the simple case without latent-overlapping groups. In this case, following Bondell and Reich (2009), note that the expanded design can be expressed as $\tilde{\bm{T}}^{\dagger}=\bm{T}\tilde{\bm{M}}^{\dagger}$ where $\tilde{\bm{M}}^{\top}=[\bm{I},\bm{D}_{\mathrm{Main}}^{\top},\bm{D}_{\mathrm{Int}}^{\top}]$ and $\tilde{\bm{M}}^{\dagger}$ is a left-inverse of $\tilde{\bm{M}}$. The latent-overlapping group formulation is a simple extension; we copy the columns of $\tilde{\bm{T}}^{\dagger}$ that correspond to $\bm{\delta}_{\mathrm{Main}}$ and append them to get $\bm{T}_{\mathrm{LOG}}$.
With this new design and parameterization in hand, we can again use the above results on projecting out the linear constraints to turn the problem into inference on an unconstrained vector $\bm{\beta}_{k}$ with a set of positive semi-definite constraints $\{\bm{F}_{g}\}_{g=1}^{2G}$ and inference proceeds identically to before.
Supplementary Material I Estimators
Here we provide further details on the estimators. In particular, we discuss estimation of Average Marginal Component Effects (AMCEs) and Average Marginal Interaction Effects (AMIEs) based on our model. We consider a traditional factorial design, where each unit receives one treatment (profile), and a conjoint design in which each unit compares two treatments (profiles). We also discuss the impact of randomization restrictions on estimators and implied changes in interpretation of estimands.
I.1 Factorial designs
I.1.1 Without restrictions on randomization
For a unit in group $k$ we have
$\Pr(Y_{i}=1\mid\bm{T}_{i},\bm{X}_{i})=\zeta_{k}(\bm{T}_{i})$ (A1)
where $i=1,2,\ldots,N$ and for $k=1,2,\ldots,K$,
$\zeta_{k}(\bm{T}_{i})\ =\ \frac{\exp(\psi_{k}(\bm{T}_{i}))}{1+\exp(\psi_{k}(\bm{T}_{i}))}.$ (A2)
We model $\psi_{k}(\bm{T}_{i})$ as
$\psi_{k}(\bm{T}_{i})\ =\ \mu+\sum_{j=1}^{J}\sum_{l=0}^{L_{j}-1}\mathbf{1}\{T_{ij}=l\}\beta^{j}_{kl}+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\sum_{l=0}^{L_{j}-1}\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}\beta^{jj^{\prime}}_{kll^{\prime}},$ (A3)
for each $k=1,2,\ldots,K$, with constraints
[…]
where $\bm{\beta}_{k}$ is a stacked column vector containing all coefficients for group $k$.
We can rewrite this to aid in the interpretation of $\bm{\beta}_{k}$ as follows:
$\text{logit}(\zeta_{k}(\bm{T}_{i}))=\mu+\sum_{j=1}^{J}\sum_{l=0}^{L_{j}-1}\mathbf{1}\{T_{ij}=l\}\beta^{j}_{kl}+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\sum_{l=0}^{L_{j}-1}\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\mathbf{1}\{T_{ij}=l,T_{ij^{\prime}}=l^{\prime}\}\beta^{jj^{\prime}}_{kll^{\prime}}.$
Thus, $\beta^{j}_{kl}-\beta^{j}_{kf}$ is the AMCE going from level $f$ to level $l$ of factor $j$ on the logit probability of $Y_{i}=1$ scale.
Let $\bm{t}$ be some combination of the $J$ factors, where $\bm{t}_{j}$ is the $j$th factor’s level and $\bm{t}_{-j}$ is the levels for all factors except $j$. This allows us to easily write, taking expectation over units in group $k$,
$\mathbb{E}\left(Y_{i}\mid Z_{i}=k,{T}_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}\right)$ $=\Pr\left(Y_{i}=1|Z_{i}=k,\bm{T}_{i,j}=l,\bm{T}_{i,-j}=\bm{t}_{-j}\right)$
$=\frac{\exp(\zeta_{k}(T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}))}{1+\exp(\zeta_{k}(T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}))},$
where $T_{ij}=l$ indicates for unit $i$ forcing factor $j$ to be assigned level $l$ and $\bm{T}_{i,-j}=\bm{t}_{-j}$ indicates forcing the assignment on all factors except for $j$ to be assigned levels as in $\bm{t}_{-j}$.
The causal effects of interest (on the original $Y$ scale) are defined as contrasts of these expectations. Without additional weighting (i.e., using traditional uniform weights for marginalization), the AMCE for level $l$ vs $f$ of factor $j$ in group $k$ is,
$\delta^{*}_{jk}(l,f)=$ $\frac{1}{M}\sum_{\bm{t_{-j}}}\mathbb{E}\left(Y_{i}\mid Z_{i}=k,{T}_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}\right)-\mathbb{E}\left(Y_{i}\mid Z_{i}=k,{T}_{ij}=f,\bm{T}_{i,-j}=\bm{t}_{-j}\right)$
$=$ $\frac{1}{M}\sum_{\bm{t_{-j}}}\frac{\exp(\zeta_{k}(T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}))}{1+\exp(\zeta_{k}(T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}))}-\frac{\exp(\zeta_{k}(T_{ij}=f,\bm{T}_{i,-j}=\bm{t}_{-j}))}{1+\exp(\zeta_{k}(T_{ij}=f,\bm{T}_{i,-j}=\bm{t}_{-j}))},$
where $M$ is the number of possible combinations of the other $J-1$ factors (e.g., if we had $J$ 2-level factors, $M=2^{J-1}$). We can estimate this by plugging in the coefficients directly. Note that, because of the nonlinear nature of the estimator, this approach is consistent (under model assumptions) but not unbiased.
Alternatively, instead of summing over all possible $\bm{t}_{-j}$, we can use the empirical distribution of $\bm{t}_{-j}$ in the sample. This potentially changes the estimand. Define estimators
$\widehat{\psi}_{k}(\bm{t})\ =\ \mu+\sum_{j=1}^{J}\sum_{l=0}^{L_{j}-1}\mathbf{1}\{t_{j}=l\}\widehat{\beta}^{j}_{kl}+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\sum_{l=0}^{L_{j}-1}\sum_{l^{\prime}=0}^{L_{j^{\prime}}-1}\mathbf{1}\{t_{j}=l,t_{j^{\prime}}=l^{\prime}\}\widehat{\beta}^{jj^{\prime}}_{kll^{\prime}}$
and
$\widehat{y}_{k}(\bm{t})=\frac{\exp(\widehat{\psi}_{k}(\bm{t}))}{1+\exp(\widehat{\psi}_{k}(\bm{t}))}$
Then we can use the following overall estimator for the AMCE:
$\frac{1}{N}\sum_{b=1}^{N}\left(\widehat{Y}_{k}(T_{bj}=l,\bm{T}_{b,-j})-\widehat{Y}_{k}(T_{bj}=f,\bm{T}_{b,-j})\right).$
This is a consistent estimator (under model assumptions) of
$\frac{1}{N}\sum_{b=1}^{N}\mathbb{E}\left(Y_{i}\mid Z_{i}=k,\bm{T}_{ij}=l,\bm{T}_{i,-j}=\bm{T}_{b,-j}\right)-\mathbb{E}\left(Y_{i}\mid Z_{i}=k,\bm{T}_{ij}=f,\bm{T}_{i,-j}=\bm{T}_{b,-j}\right)$
$=$ $\frac{1}{N}\sum_{b=1}^{N}\frac{\exp(\psi_{k}(T_{bj}=l,\bm{T}_{b,-j}))}{1+\exp(\psi_{k}(T_{bj}=l,\bm{T}_{b,-j}))}-\frac{\exp(\psi_{k}(T_{bj}=f,\bm{T}_{b,-j}))}{1+\exp(\psi_{k}(T_{bj}=f,\bm{T}_{b,-j}))},$
conditioning on the treatments we actually observed.
Now, we turn to examination of the AMIEs. Without additional weighting (i.e., using traditional uniform weights for marginalization), the AMIE for level $l$ of factor $j$ and level $q$ of factor $s$ vs $f$ of factor $j$ and level $r$ of factor $s$ in group $k$ is
$\text{AMIE}^{*}_{jsk}(l,f,q,r)=\text{ACE}^{*}(l,f,q,r)-\delta^{*}_{jk}(l,f)-\delta^{*}_{sk}(q,r)$
where
$\text{ACE}^{*}(l,f,q,r)$
$=$ $\frac{1}{M^{*}}\sum_{\bm{t_{-(j,s)}}}\mathbb{E}\left(Y_{i}\mid Z_{i}=k,T_{ij}=l,T_{is}=q,\bm{T}_{i,-(j,s)}=\bm{t}_{-(j,s)}\right)-\mathbb{E}\left(Y_{i}\mid Z_{i}=k,T_{ij}=f,T_{is}=r,\bm{T}_{i,-(j,s)}=\bm{t}_{-(j,s)}\right)$
$=$ $\frac{1}{M^{*}}\sum_{\bm{t_{-(j,s)}}}\frac{\exp(\psi_{k}(T_{ij}=l,T_{is}=q,\bm{T}_{i,-(j,s)}=\bm{t}_{-(j,s)}))}{1+\exp(\psi_{k}(T_{ij}=l,T_{is}=q,\bm{T}_{i,-(j,s)}=\bm{t}_{-(j,s)}))}-\frac{\exp(\psi_{k}(T_{ij}=f,T_{is}=r,\bm{T}_{i,-(j,s)}=\bm{t}_{-(j,s})))}{1+\exp(\psi_{k}(T_{ij}=f,T_{is}=r,\bm{T}_{i,-(j,s)}=\bm{t}_{-(j,s)}))},$
where $M^{*}$ is the number of possible combinations of the other $J-2$ factors (e.g., if we had $J$ two-level factors, $M^{*}=2^{J-2}$).
We can use the following overall estimator for the ACE:
$\widehat{\text{ACE}}^{*}(l,f,q,r)=\frac{1}{N}$ $\sum_{b=1}^{N}\widehat{Y}_{k}(T_{bj}=l,T_{bs}=q,\bm{T}_{b,-(j,s)})-\widehat{Y}_{k}(T_{bj}=f,T_{bs}=r,\bm{T}_{b,-(j,s)}).$
This is then combined with the estimators for the AMCEs to get
$\widehat{\text{AMIE}}^{*}_{jsk}(l,f,q,r)=\widehat{\text{ACE}}^{*}(l,f,q,r)-\widehat{\delta}^{*}_{jk}(l,f)-\widehat{\delta}^{*}_{sk}(q,r).$
I.1.2 With restrictions on randomization
[…]
When estimating the AMCE for level $l$ vs $f$ of factor $J-1$ in group $k$, using the model rather than the empirical distribution, we consider,
$\frac{1}{M_{def(j,h)}}\sum_{\bm{t_{-j}}:\bm{t}_{h}\notin\mathcal{S}(h,j)}\mathbb{E}\left(Y_{i}\mid Z_{i}=k,T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}\right)-\mathbb{E}\left(Y_{i}\mid Z_{i}=k,T_{ij}=f,\bm{T}_{i,-j}=\bm{t}_{-j}\right)$
$=$ $\frac{1}{M_{def(j,h)}}\sum_{\bm{t_{-j}}:\bm{t}_{h}\notin\mathcal{S}(h,j)}\frac{\exp(\psi_{k}(T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}))}{1+\exp(\psi_{k}(T_{ij}=l,\bm{T}_{i,-j}=\bm{t}_{-j}))}-\frac{\exp(\psi_{k}(T_{ij}=f,\bm{T}_{i,-j}=\bm{t}_{-j}))}{1+\exp(\psi_{k}(T_{ij}=f,\bm{T}_{i,-j}=\bm{t}_{-j}))},$
where $M_{def(j,h)}$ is the number of possible combinations of the other factors, restricted such that $\bm{t}_{h}\notin\mathcal{S}(h,j)$ (e.g., if we had $J$ 3-level factors, and some of the levels of factor $j$ were not defined for one level of factor $h$, this would be $2\times 3^{J-2}$).
To use empirical distribution, we need a way to deal with profiles that are not well defined. We can accomplish this by only aggregating over those profiles that are sensible for all levels of factor $j$. That is, we use the following estimator,
$\frac{1}{\sum_{i=1}^{N}\mathbb{I}\{T_{ih}\notin\mathcal{S}(h,j)\}}\sum_{b=1}^{N}\mathbb{I}\{T_{bh}\notin\mathcal{S}(h,j)\}\left(\widehat{Y}_{k}(T_{bj}=l,\bm{T}_{b,-j})-\widehat{Y}_{k}(T_{bj}=f,\bm{T}_{b,-j})\right).$
Consider the case where we are estimating the AMCE for “doctor” vs “gardener” for profession. Because of the randomization restriction between certain professions and level of education, we will remove any profiles that have “4th grade” as level of education. Although “gardener” with “4th grade” education is allowable under the randomization, we must remove such profiles to have an “apples-to-apples” comparison with profession of doctor, which is not allowed to have “4th grade” education. Note that we do this dropping of profiles even if we are comparing “waiter” vs “gardener” for profession, which are both allowed to have “4th grade” as level of education, to ensure that all AMCEs for profession comparable.
Similarly for the AMIEs, we restrict the profiles we marginalize over to be only those that are defined for both factors in the interactions. Let factor $j$ be restricted by some other factor $h$ and let factor $s$ be restricted by some other factor $w$. Then we have the following estimator,
$\widehat{\text{ACE}}^{*}(l,f,q,r)$
$=$ $\sum_{b=1}^{N}\frac{\mathbb{I}\{T_{bh}\notin\mathcal{S}(h,j),T_{bw}\notin\mathcal{S}(w,s)\}}{\sum_{i=1}^{N}\mathbb{I}\{T_{ih}\notin\mathcal{S}(h,j),T_{iw}\notin\mathcal{S}(w,s)\}}\Big{(}\widehat{Y}_{k}(T_{bj}=l,T_{bs}=q,\bm{T}_{b,-(j,s)})-\widehat{Y}_{k}(T_{bj}=f,T_{bs}=r,\bm{T}_{b,-(j,s)})\Big{)}.$
The relevant AMCEs should be similarly restricted within the AMIE estimator, with restrictions applied based on the restrictions for all levels both factors in the interaction.
I.2 Conjoint designs
I.2.1 Without restrictions on randomization
Consider a conjoint experiment in which each unit $i$ only compares two profiles. The response $Y_{i}$ indicates a choice between two profiles. Let $\bm{T}_{i}^{L}$ be the levels for the left profile and $\bm{T}_{i}^{R}$ be the levels for the right profile that unit $i$ sees. Here, we modify how we model $\psi_{k}$ to
$\begin{split}\psi_{k}(\bm{T}_{i}^{L},\bm{T}_{i}^{R})\ =\ &\mu+\sum_{j=1}^{J}\sum_{l\in L_{j}}\beta^{j}_{kl}\left(\mathbf{1}\left\{T^{L}_{ij}=l\right\}-\mathbf{1}\left\{T^{R}_{ij}=l\right\}\right)\\ &+\sum_{j=1}^{J-1}\sum_{j^{\prime}>j}\sum_{l\in L_{j}}\sum_{l^{\prime}\in L_{j^{\prime}}}\beta^{jj^{\prime}}_{kll^{\prime}}\left(\mathbf{1}\left\{T^{L}_{ij}=l,T^{L}_{ij^{\prime}}=l^{\prime}\right\}-\mathbf{1}\left\{T^{R}_{ij}=l,T^{R}_{ij^{\prime}}=l^{\prime}\right\}\right).\end{split}$
If we use $Y_{i}=1$ to indicate that unit $i$ picks the left profile, then we have,
$\mathbb{E}\left(Y_{i}\mid Z_{i}=k,\bm{T}_{i}^{L}=\bm{t}^{L},\bm{T}_{i}^{R}=\bm{t}^{R}\right)$ $=\Pr\left(Y_{i}=1\mid Z_{i}=k,\bm{T}_{i}^{L}=\bm{t}^{L},\bm{T}_{i}^{R}=\bm{t}^{R}\right)$
$=\frac{\exp(\psi_{k}(\bm{T}_{i}^{L}=\bm{t}^{L},\bm{T}_{i}^{R}=\bm{t}^{R}))}{1+\exp(\psi_{k}(\bm{T}_{i}^{L}=\bm{t}^{L},\bm{T}_{i}^{R}=\bm{t}^{R}))}.$
We can use the symmetry assumption that choice order does not affect the appeal of individual attributes. That is, there may be some overall preference for left or right accounted for by $\mu$, but this preference is not affected by profile attributes. Then, we can define our effects, on the original $Y$ scale, as contrasts of these expectations. Without additional weighting, the AMCE for level $l$ vs $l^{\prime}$ of factor $j$ in group $k$ is,
$\delta_{jk}(l,l^{\prime})$ $\ =\ \frac{1}{2}\mathbb{E}\left[\left\{\Pr\left(Y_{i}=1\mid Z_{i}=k,T^{L}_{ij}=l,\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i}\right)-\Pr\left(Y_{i}=1\mid Z_{i}=k,T^{L}_{ij}=l^{\prime},\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i}\right)\right\}\right.$
+$\hskip 54.2025pt\left.+\left\{\Pr\left(Y_{i}=0\mid Z_{i}=k,T^{R}_{ij}=l,\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i}\right)-\Pr\left(Y_{i}=0\mid Z_{i}=k,T^{R}_{ij}=l^{\prime},\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i}\right)\right\}\right].$
To save space, the outer expectation is over the random assignment, which corresponds to the expectation over the $\tilde{M}$ possible combinations of the two profiles on the other $J-1$ factors (e.g., if we had $J$ two-level factors, this would be $4^{J-1}$). We can again estimate this by plugging in our coefficient estimates directly.
Alternatively, instead of summing over all possible $\bm{t}^{L}_{-j}$ and $\bm{t}^{R}_{-j}$, we can use the empirical distribution of $\bm{t}^{L}_{-j}$ and $\bm{t}^{R}_{-j}$ in the sample. Define
$\widehat{Y}_{k}(\bm{t}^{L},\bm{t}^{R})=$ $\frac{\exp(\widehat{\psi}(\bm{t}^{L},\bm{t}^{R}))}{1+\exp(\widehat{\psi}(\bm{t}^{L},\bm{t}^{R}))}.$
Then we can use the estimator
$\widehat{\delta}_{jk}(l,l^{\prime})=$ $\frac{1}{2N}\sum_{i=1}^{N}\Bigg{[}\left\{\widehat{Y}_{k}(T^{L}_{ij}=l,\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i})-\widehat{Y}_{k}(T^{L}_{ij}=l^{\prime},\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i})\right\}$
$\qquad\qquad-\left\{\widehat{Y}_{k}(T^{R}_{ij}=l,\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i})-\widehat{Y}_{k}(T^{R}_{ij}=l^{\prime},\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i})\right\}\Bigg{]}.$
Now we turn to examination of the AMIEs. Without additional weighting (i.e., using traditional uniform weights for marginalization), the AMIE for level $l$ of factor $j$ and level $q$ of factor $s$ vs $m$ of factor $j$ and level $r$ of factor $s$ in group $k$ is
$\text{AMIE}_{jsk}(l,f,q,r)=\text{ACE}(l,f,q,r)-\delta_{jk}(l,f)-\delta_{sk}(q,r)$
Here we can use the estimator
$\widehat{\text{ACE}}(l,f,q,r)=$ $\frac{1}{2N}\sum_{i=1}^{N}\left[(\widehat{Y}_{k}(T^{L}_{ij}=l,T^{L}_{is}=q,\bm{T}^{L}_{i,-(j,s)},\bm{T}^{R}_{i})-\widehat{Y}_{k}(T^{L}_{ij}=f,T^{L}_{is}=r,\bm{T}^{L}_{i,-(j,s)},\bm{T}^{R}_{i})\right)$
$-\frac{1}{2N}\sum_{i=1}^{N}\left(\widehat{Y}_{k}(T^{R}_{ij}=l,T^{R}_{is}=q,\bm{T}^{R}_{i,-(j,s)},\bm{T}^{L}_{i})-\widehat{Y}_{k}(T^{R}_{ij}=f,T^{R}_{is}=r,\bm{T}^{R}_{i,-(j,s)},\bm{T}^{L}_{i})\right).$
This gives us
$\widehat{\text{AMIE}}_{jsk}(l,f,q,r)=\widehat{\text{ACE}}(l,f,q,r)-\widehat{\delta}_{jk}(l,f)-\widehat{\delta}_{sk}(q,r).$
I.2.2 With restrictions on randomization
Similar to Appendix I.1.2, adjustments to estimation need to be made when we have restricted randomizations. We again will do this by dropping profiles that have levels of factors not allowable for all levels of the factor(s) whose effects we are estimating (e.g., profiles with “4th grade” for education when estimating an effect for profession). However, now we estimate the effect for the right profile and the effect for the left profile, and then average the two (they should be equal under symmetry). When estimating the effect for the right profile, therefore, we will only drop pairings if the right profile has a level that is not allowed for some level of the factor we are estimating an effect of. For example, dropping pairings where the right profile has “4th grade” as level of education when estimating main effects of profession because “doctor” cannot have level “4th grade.” Again, this will drop more profiles than those that are not allowed under randomization to ensure an “apples-to-apples” comparison across levels of profession.
In this calculation, we use the empirical distribution for the levels of the left profile (which represents the “opponent”). Thus, the distribution of other factors for the profile we are calculating the effect of may differ than that distribution for its opponents. Similarly, when estimating the effect for the left profile, we only drop pairings in which the left profile has a restricted level for some level of the factor of interest. Estimation for the AMIE under randomization restrictions follows similarly.
Supplementary Material J Quantification of Uncertainty
We quantify uncertainty in our parameter estimates by inverting the negative Hessian of the log-posterior at the estimates $\hat{\bm{\theta}}$, i.e. $\left[-\frac{\partial}{\partial\bm{\theta}\bm{\theta}^{T}}\log p(\bm{\theta}|Y_{i})\right]_{\bm{\theta}=\hat{\bm{\theta}}}$ or $\mathcal{I}(\hat{\bm{\theta}})$. This can be stably and easily computed using terms from the AECM algorithm following Louis (1982)’s method. Specifically, consider the model from the main text augmented with $Z_{i}$, i.e. the group memberships. Recall that $z_{ik}=\mathbf{1}\{Z_{i}=k\}$ for notational simplicity.
$L^{c}(\bm{\theta})=$ $\sum_{i=1}^{N}\left[\sum_{k=1}^{K}z_{ik}\log(\pi_{ik})+z_{ik}\log L(Y_{i}\mid\bm{\beta}_{k})\right]+$ (A1)
$\sum_{k=1}^{K}m\log(\lambda)+m\gamma\log(\bar{\pi}_{k})-\lambda\bar{\pi}^{\gamma}_{k}\left[\sum_{g=1}^{G}\xi_{gk}\sqrt{\bm{\beta}^{\top}_{k}\bm{F}_{g}\bm{\beta}_{k}}\right]+\log p(\{\bm{\phi}_{k}\}).$
Louis (1982) notes that equation can be used to compute $\mathcal{I}_{L}(\hat{\bm{\theta}})$, where the subscript $L$ denotes its computation via this method.
$\mathcal{I}_{L}(\hat{\bm{\theta}})=E_{p\left(\{Z_{i}\}_{i=1}^{N}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N},\hat{\bm{\theta}}\right)}\left[-\frac{\partial L^{c}(\bm{\theta})}{\partial\bm{\theta}\bm{\theta}^{\top}}\right]-\mathrm{Var}_{p\left(\{Z_{i}\}_{i=1}^{N}\mid\{Y_{i},\bm{X}_{i},\bm{T}_{i}\}_{i=1}^{N},\hat{\bm{\theta}}\right)}\left[\frac{\partial L^{c}(\bm{\theta})}{\partial\bm{\theta}}\right]$ (A2)
To address the issue with the non-differentiability of the penalty on $\bm{\beta}$ (and thus $L^{c}(\bm{\theta})$), we follow the existing research in two ways. First, for restrictions that are sufficiently close to binding, we assume them to bind and estimate the uncertainty given those restrictions. That is, we identify the binding restrictions such that $\sqrt{\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}}$ is sufficiently small (say $10^{-4}$) and note that if these are binding, we can use the null space projection technique to transform $\bm{\beta}_{k}$ such that it lies in an unconstrained space.
To further ensure stability, we modify the penalty with a small positive constant $\epsilon\approx 10^{-4}$ to ensure that the entire objective is (twice) differentiable. For notational simplicity, we derive the results below assuming $\bm{\beta}_{k}$ represent the parameter vector after projecting into a space with no linear constraints. The approximated log-posterior is shown below and denoted with a tilde. We thus evaluate $\mathcal{I}_{L}(\hat{\bm{\theta}})$ using $\tilde{L}^{c}$ in place of $L^{c}$.
$\tilde{L}^{c}(\bm{\theta})=$ $\sum_{i=1}^{N}\left[\sum_{k=1}^{K}z_{ik}\log(\pi_{ik})+z_{ik}\log L(y_{i}|\bm{\beta}_{k})\right]+$ (A3)
$\sum_{k=1}^{K}m\log(\lambda)+m\gamma\log(\bar{\pi}_{k})-\lambda\bar{\pi}^{\gamma}_{k}\left[\sum_{g=1}^{G}\xi_{gk}\sqrt{\bm{\beta}^{\top}_{k}\bm{F}_{g}\bm{\beta}_{k}+\epsilon}\right]+\log p(\{\bm{\phi}_{k}\})$
This procedure has some pleasing properties that mirror existing results on approximate standard errors after sparse estimation; consider a simple three-level case: $\beta^{j}_{1},\beta^{j}_{2},\beta^{j}_{3}$. If $\beta^{j}_{1}$ and $\beta^{j}_{2}$ are fused, then their approximate point estimates and standard errors will be identical but crucially not zero. This is because while their difference is zero and assumed to bind with no uncertainty, this does not imply that the effects, themselves, have no uncertainty: $\beta^{j}_{1}-\beta^{j}_{2}$ will have a standard error of zero in our method. This thus mirrors the results from Fan and Li (2001) where effects that are shrunken to zero by the LASSO are not estimated with any uncertainty. One might relax this with fully Bayesian approaches in future research.
Second, note that if all levels are fused together, i.e. $\beta^{j}_{1}=\beta^{j}_{2}=\beta^{j}_{3}$, then all point estimates must be zero by the ANOVA sum-to-zero constraint and all will have an uncertainty of zero. Thus, when an entire factor is removed from the model, the approximate standard errors return a result consist with existing research.
J.1 Derivation of Hessian
To calculate the above terms, the score and gradient of $\tilde{L}^{c}$ are required. They are reported below:
$\tilde{S}^{c}(\mu)$ $=\sum_{i=1}^{N}\left[\sum_{k=1}^{K}z_{ik}(Y_{i}-p_{ik})\right]$
$\tilde{S}^{c}(\bm{\beta}_{k})$ $=\sum_{i=1}^{N}z_{ik}\cdot(Y_{i}-p_{ik})\tilde{\bm{T}}_{i}-\lambda\bar{\pi}^{\gamma}_{k}\sum_{g=1}^{G}\xi_{gk}(\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k})^{-1/2}\cdot\bm{F}_{g}\bm{\beta}_{k}$
$\begin{split}\tilde{S}^{c}(\bm{\phi}_{k})&=\sum_{i=1}^{N}\left[z_{ik}-\pi_{ik}\right]\bm{X}_{i}+\frac{\partial\log p(\{\bm{\phi}_{k}\})}{\partial\bm{\phi}_{k}}+\\ &\sum_{k^{\prime}=1}^{K}m\gamma\frac{\partial\log(\bar{\pi}_{k^{\prime}})}{\partial\bm{\phi}_{k}}-\lambda\gamma\bar{\pi}_{k^{\prime}}^{\gamma-1}\cdot\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\cdot\left[\sum_{g=1}^{G}\xi_{g,k^{\prime}}\sqrt{\bm{\beta}_{k^{\prime}}^{\top}\bm{F}_{g,k^{\prime}}\bm{\beta}_{k^{\prime}}}\right]\end{split}$
$H^{c}(\mu,\mu)$ $=\sum_{i=1}^{N}\left[-\sum_{k=1}^{K}z_{ik}p_{ik}(1-p_{ik})\right]$
$H^{c}(\mu,\bm{\beta}_{k})$ $=-\left[\sum_{i=1}^{N}z_{ik}p_{ik}(1-p_{ik})\tilde{\bm{T}}_{i}\right]$
$H^{c}(\bm{\beta}_{k},\bm{\beta}_{k})$ $=-\left[\sum_{i=1}^{N}z_{ik}\cdot p_{ik}(1-p_{ik})\tilde{\bm{T}}_{i}\tilde{\bm{T}}_{i}^{\top}\right]-\lambda\bar{\pi}^{\gamma}_{k}\sum_{g=1}^{G}\xi_{gk}\bm{D}_{gk}$
where $\left[\bm{D}_{gk}\right]_{a,b}=-\left(\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}\right)^{-3/2}\bm{\beta}_{k}^{\top}\left[\bm{F}_{g}\right]_{a}\bm{\beta}_{k}^{\top}\left[\bm{F}_{g}\right]_{b}+\left(\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k}\right)^{-1/2}\left[\bm{F}_{g}\right]_{a,b}$.
$H^{c}(\left[\bm{\beta}_{k}\right]_{i},\bm{\phi}_{\ell})$ $=-\lambda\gamma\bar{\pi}^{\gamma-1}_{k}\left[\sum_{g=1}^{G}\xi_{gk}(\bm{\beta}_{k}^{\top}\bm{F}_{g}\bm{\beta}_{k})^{-1/2}\cdot\bm{\beta}_{k}^{\top}\left[\bm{F}_{g}\right]_{i}\right]\frac{\partial\bar{\pi}_{k}}{\partial\bm{\phi}_{\ell}}$
$H^{c}(\bm{\phi}_{k},\bm{\phi}_{\ell})$ $=\sum_{i=1}^{N}-\left[\left(I[k=\ell]-\pi_{ik}\right)\pi_{i\ell}\right]\bm{X}_{i}\bm{X}_{i}^{\top}+\frac{\partial^{2}\log p(\{\bm{\phi}_{k}\})}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}+\sum_{k^{\prime}=1}^{K}m\gamma\frac{\partial\log(\bar{\pi}_{k^{\prime}})}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}+$
$\sum_{k^{\prime}=1}^{K}-\lambda\gamma$ $\left[\sum_{g=1}^{G}\xi_{g,k^{\prime}}\sqrt{\bm{\beta}_{k^{\prime}}^{\top}\bm{F}_{g,k^{\prime}}\bm{\beta}_{k^{\prime}}}\right]\left[I(\gamma\notin\{0,1\})\cdot(\gamma-1)\bar{\pi}_{k^{\prime}}^{\gamma-2}\cdot{\color[rgb]{0,0,0}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\right]}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{\ell}}\right]^{\top}+\bar{\pi}_{k^{\prime}}^{\gamma-1}\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}\right]$
The above results use the following intermediate derivations:
$\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}$ $=\frac{1}{N}\sum_{i=1}^{N}\pi_{i,k^{\prime}}\left[I(k=k^{\prime})-\pi_{ik}\right]\bm{X}_{i}$
$\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}$ $=\frac{1}{N}\sum_{i=1}^{N}\left[\pi_{i,k^{\prime}}\left(I(k^{\prime}=\ell)-\pi_{i\ell}\right)\left(I(k=k^{\prime})-\pi_{ik}\right)-\pi_{i,k^{\prime}}\pi_{ik}\left(I(k=\ell)-\pi_{i\ell}\right)\right]\bm{X}_{i}\bm{X}_{i}^{\top}$
$\frac{\partial\log(\bar{\pi}_{k^{\prime}})}{\partial\bm{\phi}_{k}}$ $=\frac{1}{\bar{\pi}_{k^{\prime}}}\cdot\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}$
$\frac{\partial\log(\bar{\pi}_{k^{\prime}})}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}$ $=-\frac{1}{\bar{\pi}_{k^{\prime}}^{2}}\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}}\right]\left[\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{\ell}}\right]^{\top}+\frac{1}{\bar{\pi}_{k^{\prime}}}\cdot\frac{\partial\bar{\pi}_{k^{\prime}}}{\partial\bm{\phi}_{k}\bm{\phi}_{\ell}^{\top}}$
Second, the variance of $\tilde{S}^{c}(\bm{\theta})$ over $p(\{z_{ik}\}\mid\bm{\theta})$. This is derived blockwise below.
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\beta}_{k}),\tilde{S}^{c}(\bm{\beta}_{\ell})\right]$ $=\sum_{i=1}^{N}(Y_{i}-p_{ik})\cdot(Y_{i}-p_{i\ell})\cdot\mathbb{E}(z_{ik})\left(I(k=\ell)-\mathbb{E}(z_{i\ell})\right)\tilde{\bm{T}}_{i}\tilde{\bm{T}}_{i}^{\top}$
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\beta}_{k}),\tilde{S}^{c}(\bm{\phi}_{\ell})\right]$ $=\sum_{i=1}^{N}(Y_{i}-p_{ik})\cdot\mathbb{E}(z_{ik})\left(I(k=\ell)-\mathbb{E}(z_{i\ell})\right)\tilde{\bm{T}}_{i}\bm{X}_{i}^{\top}$
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\phi}_{k}),\tilde{S}^{c}(\bm{\phi}_{\ell})\right]$ $=\sum_{i=1}^{N}\mathbb{E}(z_{ik})\left(I(k=\ell)-\mathbb{E}(z_{i\ell})\right)\bm{X}_{i}\bm{X}_{i}^{\top}$
$\mathrm{Cov}\left[\tilde{S}^{c}(\mu),\tilde{S}^{c}(\mu)\right]$ $=\sum_{i=1}^{N}\left[\sum_{k=1}^{K}\sum_{k^{\prime}=1}^{K}\mathbb{E}(z_{ik})\left(I(k=k^{\prime})-\mathbb{E}(z_{ik^{\prime}})\right)(Y_{i}-p_{ik})(Y_{i}-p_{ik^{\prime}})\right]$
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\phi}_{k}),\tilde{S}^{c}(\mu)\right]$ $=\sum_{i=1}^{N}\left[\sum_{k^{\prime}=1}^{K}\mathbb{E}(z_{ik})\left(I(k=k^{\prime})-\mathbb{E}(z_{ik^{\prime}})\right)(Y_{i}-p_{ik^{\prime}})\bm{X}_{i}\right]$
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\beta}_{k}),\tilde{S}^{c}(\mu)\right]$ $=\sum_{i=1}^{N}\left[\sum_{k^{\prime}=1}^{K}\mathbb{E}(z_{ik})\left(I(k=k^{\prime})-\mathbb{E}(z_{ik^{\prime}})\right)(Y_{i}-p_{ik})(Y_{i}-p_{ik^{\prime}})\tilde{\bm{T}}_{i}\right]$
This provides all terms needed to compute $\mathcal{I}_{L}(\hat{\bm{\theta}})$.
J.2 Repeated Observations
Now consider the case of repeated observations per individual $i$. In this scenario, each individual $i$ performs $N_{i}$ tasks. Note, after augmentation, the score has exactly the same form and thus the complete Score $\tilde{S}^{c}$ and Hessian $\tilde{H}^{c}$ are identical where the sum merely now runs over $\sum_{i=1}^{N}\sum_{m=1}^{N_{i}}$. The average for $\bar{\pi}_{k}$ is similarly a weighted average by $N_{i}$, although note that often each respondent answers an identical number of tasks so it is, effectively, the same as before. The covariance of $\tilde{S}^{c}$ is adjusted as shown below.
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\beta}_{k}),\tilde{S}^{c}(\bm{\beta}_{\ell})\right]$ $=\sum_{i=1}^{N}\mathbb{E}(z_{ik})\left(I[k=\ell]-\mathbb{E}(z_{i\ell})\right)\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}\right]\left[\sum_{m^{\prime}=1}^{N_{i}}(Y_{im}-p_{im\ell})\tilde{\bm{T}}_{im^{\prime}}^{\top}\right]$
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\beta}_{k}),\tilde{S}^{c}(\bm{\phi}_{\ell})\right]$ $=\sum_{i=1}^{N}\mathbb{E}(z_{ik})\left(I[k=\ell]-\mathbb{E}(z_{i\ell})\right)\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}\right]\bm{X}_{i}^{\top}$
$\mathrm{Cov}\left[\tilde{S}^{c}(\mu),\tilde{S}^{c}(\mu)\right]$ $=$
$\sum_{i=1}^{N}\Bigg{[}\sum_{k=1}^{K}\sum_{k^{\prime}=1}^{K}\mathbb{E}(z_{ik})$ $\left(I[k=k^{\prime}]-\mathbb{E}(z_{ik^{\prime}})\right)\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}\right]\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}^{\top}\right]\Bigg{]}$
$\mathrm{Cov}\left[\tilde{S}^{c}(\bm{\beta}_{k}),\tilde{S}^{c}(\mu)\right]$ $=\sum_{i=1}^{N}\left[\sum_{k^{\prime}=1}^{K}\mathbb{E}(z_{ik})\left(I[k=k^{\prime}]-\mathbb{E}(z_{ik^{\prime}})\right)\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk})\tilde{\bm{T}}_{im}\right]\left[\sum_{m=1}^{N_{i}}(Y_{im}-p_{imk^{\prime}})\right]\right]$
J.3 Standard Errors on Other Quantities of Interest
Given the above results, we derive an approximate covariance matrix on $\hat{\bm{\theta}}$. We calculate uncertainty on other quantities of interest, e.g. AMCE and marginal effects, using the multivariate delta method. As almost all of our quantities of interest can be expressed as (weighted) sums or averages over individuals $i\in\{1,\cdots,N\}$, calculating the requisite gradient for the multivariate delta method simply requires calculating the relevant derivative for each observation. For example, all derivatives needed in the AMCE are of the following form; see Appendix I for more details.
$\frac{\partial}{\partial\bm{\theta}}\left[\frac{\exp(\psi_{ik})}{1+\exp(\psi_{ik})}\right]$
Supplementary Material K Simulations
We detail our simulations and provide additional results in this section.
K.1 Setup
We generate the $\bm{\beta}_{k}$ used in our simulations following Equation 3.5 and calibrating their implied AMCEs to be roughly comparable to the magnitude found in our empirical example, i.e. ranging between around $-0.30$ and $0.30$. The $\bm{\beta}_{k}$ and $\{\bm{\phi}_{k}\}_{k=2}^{3}$ used in all simulations are determined using one draw from the following procedure:
  • Simulating $\bm{\beta}_{k}$:
    1. 1.
      For each factor $j$ and group $k$, draw the number of unique levels $u$ with equal probability from $\{1,2,3\}$.
    2. 2.
      Draw $u$ normal random variables independently from $N(0,1/3)$; call these $b^{j}_{ku}$.
    3. 3.
      For $u=1$, set $\beta^{j}_{kl}=0$
    4. 4.
      For $u=3$, de-mean $\{b^{j}_{ku}\}_{u=1}^{3}$ drawn in (2) and set all $\beta^{j}_{kl}$ equal to the corresponding value.
    5. 5.
      For $u=2$, assign $b^{j}_{k3}$ equal to one of the two $b^{j}_{ku}$ with equal probability. De-mean the $\{b^{j}_{ku}\}_{u=1}^{3}$ and set $\beta^{j}_{kl}$ equal to the corresponding values.
  • Simulating $\bm{\phi}_{k}$: $\{\bm{\phi}_{k}\}_{k=1}^{K}\sim N(\bm{0},2\cdot\bm{I})$
To evaluate our method, we calculate the AMCEs in each group simulations using Monte Carlo simulation where we sample 1,000,000 pairs of treatment profiles for the other attributes to marginalize over the other factors. The distribution of the $\bm{\beta}_{k}$ and average marginal component effects (with a baseline level of ‘1’) used in the simulations are shown below:
Figure A3: The distribution of parameters and AMCEs used in the simulation.
For each simulation, we draw $N$ individuals who rate $T$ profiles where $(N,T)\in\{(1000,5),(2000,10)\}$. For each individual $i$, we draw its moderators $\bm{x}_{i}$ from a correlated multivariate normal where $\bm{x}_{i}\sim N(\bm{0}_{5},\bm{\Sigma})$ with $\bm{\Sigma}_{ij}=0.25^{|i-j|}$ for $i,j\in\{1,\cdots,5\}$. The distribution of group assignment probabilities $\pi_{ik}$ is shown below from one million Monte Carlo simulation draws of $[1,\bm{x}_{i}^{\top}]$.
Figure A4: Group Membership Probabilities
We see that the members are well-separated; the groups are somewhat unbalanced, i.e. $\bar{\bm{\pi}}=[0.217,0.261,0.522]$. If we consider the maximum probability for each person $i$, i.e. $\pi^{*}_{i}=\max_{k\in\{1,2,3\}}\pi_{ik}$, this distribution has a median of 0.93, a 25th percentile of 0.75 and a 75th percentile of 0.99.
In terms of simulating the treatment profiles and outcome, for each individual $i$, we draw a group membership $Z_{i}$ using $\bm{\pi}_{i}$ generating using $\bm{X}_{i}$, $\bm{\phi}$ and Equation 3.5. For each task $t$, we then randomly draw a pair of treatments and then, given $Z_{i}$, draw the outcome $Y_{i}$ given their observed treatments using the model in the main text.
After estimating our model with $K=3$, we resolve the problem of label switching by permuting our estimate group labels to minimize the absolute error between the estimated posterior membership probabilities $\{E[z_{ik}|\bm{\theta}]\}_{k=1}^{K}$ and $\bm{z}_{i}$ (the one-hot assignment of group membership).
K.2 Additional Results
We provide additional simulation results to complement those presented in the main text. Figure A5 presents the results for the simulations in the main text when considering the $\bm{\beta}_{k}$ (instead of the AMCE). It shows a similar pattern of some bias even at the larger sample size.
Figure A5: The empirical performance of the proposed estimator on simulated data. The black squares indicate the effects estimated in each group with the smaller sample size (1,000 people completing 5 tasks); the red crosses indicate effects estimated with the larger sample size (2,000 people completing 10 tasks).
To address this issue, we consider an alternative procedure based on sample splitting. We fit the model using half of the data (selected at random) and then refit the model. To refit the model, we hold fixed the sparsity pattern estimated in the original estimation hold (i.e., which levels are fused together) using a tolerance of $10^{-3}$. We also fix the estimated moderator relationship, i.e. $\pi_{k}(\bm{X}_{i})$, and only estimate the treatment effect coefficients after fusion. Algorithm A2 states the procedure. To calculate the average marginal effects, as noted in Appendix I, we use the empirical distribution of treatments to marginalize over other factors. In this split version, we also use the distribution from the full dataset.
Algorithm A2 Refitting Procedure
1. Randomly split the observations $i\in\{1,\cdots,N\}$ into two groups indexed by $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$
2. Using the data $i\in\mathcal{I}_{1}$, estimate the parameters of the model using Algorithm A1 in the main text. Define the resulting parameters from this as $\tilde{\bm{\theta}}$: $\{\tilde{\bm{\beta}}_{k}\}_{k=1}^{K}$, $\{\tilde{\bm{\phi}}_{k}\}_{k=2}^{K},~{}\tilde{\mu}$
3. Fuse levels $l$ and $l^{\prime}$ of factor $j$ for group $k$ where the following condition holds for tolerance $\epsilon$
$\max\left\{\left|\tilde{\beta}^{j}_{kl}-\tilde{\beta}^{j}_{kl^{\prime}}\right|\right\}\bigcup\left\{\bigcup_{j^{\prime}\neq j}\bigcup_{m=0}^{L_{j^{\prime}}-1}\left|\tilde{\beta}^{jj^{\prime}}_{klm}-\tilde{\beta}^{jj^{\prime}}_{kl^{\prime}m}\right|\right\}\leq\epsilon$
For each combination where this is satisfied, construct matrices $\bm{R}_{k}$ that contain the required equality constraints, i.e. where $\bm{R}_{k}^{T}\tilde{\bm{\beta}}_{k}$ ensures that $\tilde{\beta}^{j}_{kl}=\tilde{\beta}^{j}_{kl^{\prime}}=0$ and/or $\tilde{\beta}^{jj^{\prime}}_{klm}-\tilde{\beta}^{jj^{\prime}}_{kl^{\prime}m}=0$. Define $\tilde{\pi}_{k}(\bm{X}_{i})$ as follows:
$\tilde{\pi}_{k}(\bm{X}_{i})=\frac{\exp(\bm{X}_{i}^{\top}\tilde{\bm{\phi}}_{k})}{\sum_{k^{\prime}=1}^{K}\exp(\bm{X}_{i}^{\top}\tilde{\bm{\phi}}_{k^{\prime}})}$
4. Using the other half of the data $i\in\mathcal{I}_{2}$, estimate the refit parameters for the treatment effects, where $\bm{C}$ contains the original sum-to-zero constraints discussed in the main text.
$\{\hat{\bm{\beta}}^{\mathrm{refit}}_{k}\}_{k=1}^{K},\hat{\mu}^{\mathrm{refit}}=\operatornamewithlimits{argmax}_{\{\bm{\beta}_{k}\}_{k=1}^{K},~{}\mu}~{}\sum_{i\in\mathcal{I}_{2}}\log\left(\sum_{k=1}^{K}\tilde{\pi}_{k}(\bm{X}_{i})\zeta_{k}(\bm{T}_{i})^{Y_{i}}\{1-\zeta_{k}(\bm{T}_{i})\}^{1-Y_{i}}\right)\quad\mathrm{s.t.}\quad\bm{C}^{T}\bm{\beta}_{k}=\bm{0},~{}\bm{R}_{k}^{T}\bm{\beta}_{k}=\bm{0}$
Figure A6 compares the estimators from the split sample and full data (“Full Sample”, i.e. the methods shown in the main text) approaches. It shows the distribution of the root mean-squared error (RMSE), bias, and coverage across the estimated AMCE and coefficients. We split the results by whether the true underlying effect is zero to compare differences across those cases. We also consider one even larger sample size (4,000 respondents with 10 tasks) to examine a scenario where the split sample method has the same amount of data as the full sample method for the second step in the estimation process.
(a) Results for AMCE
(b) Results for $\bm{\beta}_{k}$
Figure A6: The distribution of performance for each estimator across sample sizes. The top figure shows results for the AMCE; the lower figure shows results for the coefficients $\bm{\beta}_{k}$. Inside each figure, results are split by whether the true effect is zero (“Zero True Effect”) or not (“Non-Zero True Effect”). The boxplot shows the distribution across all effects for each group. For the plots on RMSE and bias, the solid vertical line indicates zero. For coverage, the solid line indicates 95% coverage and the dashed line indicates 90%.
The figure corroborates the initial results. Specifically, the full data method has non-trivial bias that decreases slowly even at the largest sample sizes. By contrast, the bias is small in the split sample method. As the panel on coverage shows, this results in considerably better coverage—especially for quantities with a non-zero true effect. At the two larger sample sizes, the median frequentist coverage of the split sample method is close to the nominal 95%, with a few outliers that have low coverage. In terms of RMSE, the methods perform similarly.
K.3 Robustness to Misspecification
As noted in the main text, our methodology is not predicated on the assumption that the true data generating process is a mixture model. Rather, fitting a mixture model or a mixture of experts model is equivalent to finding maximally heterogeneous groups. Nevertheless, we consider a simulation setting in which the true data generating process is a mixture model. Under this assumption, we explore how the specification of different parts of the model (e.g., $K$ and the choice of moderators) affects performance. Specifically, we explore different choices of $K$ and misspecification of the moderator model $\pi_{k}(\bm{X}_{i})$ from the ones used to generate the data.
K.3.1 Data-Driven Choice of $K$
First, as noted in the main text, a common approach to choosing $K$ can be information criterion. We use the BIC to calibrate our choice of $\lambda$, i.e. pick the $\lambda$ that minimizes the BIC. In our simulations, we compare the BIC across $K\in\{1,2,3,4\}$ to see which it would suggest choosing. Table K.3.1 reports the probability of each $K$ being chosen across 1,000 simulations. It shows that, even for the smallest data size, the BIC correctly identifies $K=3$. The probability of correct selection rises as the sample size grows. However, as we note in the main text, this simulation example has relatively well separated clusters, and correctly specified likelihoods, and thus the information criterion approach is expected to perform well.
Sample Size $K=1$ $K=2$ $K=3$ $K=4$
1,000 People (5 Tasks) 0 0.01 0.941 0.049
2,000 People (10 Tasks) 0 0.00 0.999 0.001
4,000 People (10 Tasks) 0 0.00 0.994 0.006
Table A2: Probability of $K$ being chosen using smallest BIC
Other criterion based on cross-validation—e.g., splitting the sample and taking the model with the highest out-of-sample predictive likelihood or lowest RMSE—also show a high probability of choosing $K=3$ (84% for the smallest sample size and 97-98% for the larger sample sizes).
K.3.2 Effect of Choice of $K$ on Estimates
We first consider how different choices of $K$ impact our results in the simulation study. To do this, we focus on the CAMCE discussed in the main text (Section 5.4) as this quantity is comparable across models with different $K$. For each individual $i$, we calculate our estimate of CAMCE using their moderators $\bm{X}_{i}$ and compare this against the true value, which can be calculated by plugging in the true values of $\pi_{k}(\bm{X}_{i})$ and $\delta_{jk}(l,l^{\prime})$ into Equation (5.2). We run models with $K\in\{2,3,4\}$ with both split-sample and full data methods discussed above.
Figure A7 shows a binned scatterplot of the true CAMCEs against the estimated CAMCEs for each individual $i$, i.e., for all true CAMCE in a bin, what is the average estimated CAMCE? As above, it shows that for the correct choice of $K=3$, the estimates track the truth well. Interestingly, $K=4$ also shows good performance but $K=2$ shows some weaker performance, especially for certain ranges of the true CAMCE.
Figure A7: The binned scatterplot of the true CAMCEs versus the estimated CAMCEs. Results are shown for different sample sizes and estimation method (e.g., full data versus split sample). The color of the dot indicates the number of groups $K$.
We also compute the marginalized error (i.e., the error in the estimated CAMCE vs the true CAMCE, averaged across all people and CAMCEs estimated in a simulation) and RMSE of the estimated CAMCEs. Figure A8 plots the distribution of RMSE and marginalized error across the 1000 simulations. Consistent with our earlier results, the figure shows that the full sample method for all choice of $K$ has some non-vanishing bias while the split-sample method exhibits a considerably smaller error. Further, while the estimated error looks similar for $K\in\{2,3,4\}$, the correct choice ($K=3$) has lower RMSE than either $K=2$ or $K=4$. The results for $K=4$ are comparable to those for $K=3$, but the case of $K=2$ sees a considerably worse performance.
Figure A8: The distribution of performance across simulations. The top panel shows the performance in terms of RMSE and marginalized error, across all individuals and CAMCEs, for the model fit on the entire dataset. The bottom panel shows the results for a method estimated using the split sample method. The color of the boxplot indicates the number of groups $K$.
Next, we consider how different choices of $K$ affect the ability to recover the average marginal effect. To do this, we average the CAMCE across all individuals used to fit the model and compare that AMCE in the population. Figure A9 plots the bias of the estimated AMCE by aggregating the individual-level effects; it is largely unaffected by the choice of $K$, corroborating Figure A8. As expected, there is regularization bias for the full data method that using the split sample approach eliminates.
Figure A9: The distribution of bias in AMCEs by averaging CAMCEs by different $K$
As a final illustration on the choice of $K$, we also examine how much variability in the true CAMCE is explained by the estimated groups, inspired by how one might assess the quality of clustering in $k$-means. We compute this as follows: For each observation $i$, obtain its estimated group membership probabilities $\hat{\pi}_{k}(\bm{X}_{i})$ for $k\in\{1,\cdots,K\}$. Using its true CAMCE, i.e. $\mathrm{CAMCE}^{*}_{j}(l,l^{\prime};\bm{X}_{i})$, compute the total variability in CAMCE across the $N$ units and the between-group variability using $\hat{\pi}_{k}$ as group weights. Formally, we compute $B_{K}$ and the total variability $T$.
$B_{K}$ $=\sum_{k=1}^{K}\sum_{j=1}^{J}\sum_{l^{\prime}_{j}=1}^{L_{j}-1}N_{k}\left[\overline{\mathrm{CAMCE}}^{*}_{k,j}(l_{j},l^{\prime}_{j})-\overline{\mathrm{CAMCE}}^{*}_{j}(l_{j},l^{\prime}_{j})\right]^{2};\quad N_{k}=\sum_{i=1}^{N}\hat{\pi}_{k}(\bm{X}_{i});$
$T$ $=\sum_{j=1}^{J}\sum_{l^{\prime}_{j}=1}^{L_{j}-1}\sum_{i=1}^{N}\left[\mathrm{CAMCE}^{*}_{j}(l_{j},l^{\prime}_{j};\bm{X}_{i})-\overline{\mathrm{CAMCE}}^{*}_{j}(l_{j},l^{\prime}_{j})\right]^{2}$
$\overline{\mathrm{CAMCE}}^{*}_{k,j}$ $=\frac{1}{N_{k}}\sum_{i=1}^{N}\hat{\pi}_{k}(\bm{X}_{i})\cdot\mathrm{CAMCE}^{*}_{j}(l_{j},l_{j}^{\prime};\bm{X}_{i});\quad\overline{\mathrm{CAMCE}}^{*}_{j}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{CAMCE}^{*}_{j}(l_{j},l_{j}^{\prime};\bm{X}_{i})$
Figure A10 reports the ratio of the between-group variability over the total variability across the 1,000 simulations for $K\in\{2,3,4\}$. With $K=2$, we already able to explain around 50% of the variability in the data. As expected, $K=2$ shows considerably lower $B_{K}/T$ than higher $K$’s, suggesting its groups are less distinct—or, equivalently, more internally heterogeneous—than $K\in\{3,4\}$. There is limited improvement in quality with $K=4$, which is consistent with the earlier results that the correct choice ($K=3$) adequately summarizes the variability in the data.
Figure A10: The distribution of $B_{K}/T$ across simulations. The top panel shows results for the model fit on the entire dataset. The bottom panel shows the results for a method estimated using the split sample method.
K.3.3 Misspecified Moderators
We next consider how misspecifying the model for the moderators $\pi_{k}(\bm{X}_{i})$ affects our simulated results. We show this in two ways; first, we fit a model with no moderators, that is, $\bm{X}_{i}=1$. While this model has a number of limitations—e.g., for classifying and predicting heterogeneous effects for new individuals, it is a useful benchmark. Second, instead of using the true moderators (e.g., $\bm{X}_{i}$), we assume the researcher only has available the following non-linear transformations of the moderators (following Kang and Schafer 2007) and uses those instead:
$\bm{A}_{i,1}$ $=\sqrt{3}\exp(\bm{X}_{i,1}/2)-2$
$\bm{A}_{i,2}$ $=\sqrt{3}\bm{X}_{i,2}/\left[1+\exp(\bm{X}_{i,1})\right]$
$\bm{A}_{i,3}$ $=1/19\left[\bm{X}_{i,1}+\bm{X}_{i,3}+0.6\right]^{3}$
$\bm{A}_{i,4}$ $=1/3\left[\bm{X}_{i,2}+\bm{X}_{i,4}\right]^{2}-1$
$\bm{A}_{i,5}$ $=2.5\sqrt{|\bm{X}_{i,5}+\bm{X}_{i,1}|}-2.5.$
We rescale the moderators $\{\bm{A}_{i}\}_{i=1}^{N}$ to have zero mean and unit variance in each simulated dataset.
Figure A11 replicates Figure A6 on the performance on estimating the AMCE where we show results with all moderators (i.e., in Figure A6) and with both types of mis-specification (“No Moderators” and “Non-Linear Transf.” when $\bm{A}_{i}$ are used).
Figure A11: The distribution of performance for each estimator across sample sizes, with and without moderators. Inside each figure, results are split by whether the true effect is zero (“Zero True Effect”) or not (“Non-Zero True Effect”). The boxplot shows the distribution across all effects for each group. For the plots on RMSE and bias, the solid vertical line indicates zero. For coverage, the solid line indicates 95% coverage and the dashed line indicates 90%.
It shows that, for the smallest sample size, the no-moderator model incurs a penalty in terms of the RMSE of the estimated AMCEs, although it does not have considerably larger bias. At larger sample sizes, the difference between the moderator and no-moderator models decreases. With moderators that are included but mis-specified using some non-linear transformation, the performance is rather close to the one that uses the correct moderators.
To further illustrate the impact of excluding moderators, Figure A12 plots the estimated average posterior and posterior predictive probability (i.e., $\hat{\pi}_{k}(\bm{X}_{i})$) in the group corresponding to the individual’s sampled $Z_{i}$ for all observations in the estimation data. It shows, as expected, that using the correctly specified moderators results in a considerably higher probability of each individual being assigned to group that corresponds to their sampled $Z_{i}$. The model with included but mis-specified moderators (“Non-Linear Transf.”) is somewhere between the model without moderators and the correctly specified one.
Figure A12: The average probability that is assigned to the group corresponding to an individual’s sampled $Z_{i}$, showing the distributions across simulations.
Supplementary Material L Additional Results for Immigration Conjoint Experiment
We provide some additional results for our main empirical analysis. First, focusing on the three-group model, we report a different quantity of interest. We use an analogue to the “marginal means” estimator in Leeper, Hobolt and Tilley (2020). We compute the probability of a profile being chosen without specifying a baseline category. The equation is shown below for the forced choice case; note it consists of two of the terms used for the AMCE.
$\mathrm{MM}_{jk}(l)$ $\ =\ \frac{1}{2}\mathbb{E}\left[\left\{\Pr\left(Y_{i}=1\mid Z_{i}=k,T^{L}_{ij}=l,\bm{T}^{L}_{i,-j},\bm{T}^{R}_{i}\right)+\Pr\left(Y_{i}=0\mid Z_{i}=k,T^{R}_{ij}=l,\bm{T}^{R}_{i,-j},\bm{T}^{L}_{i}\right)\right\}\right].$ (A1)
The below plot ignores randomization restrictions when estimating this quantity to center the estimate around 0.50 as in Leeper, Hobolt and Tilley (2020). The results are substantively similar to the analysis in shown in the main paper using AMCEs.
Figure A13: Estimated average marginal means using a three-group (right) analysis. The point estimates and 95% Bayesian credible intervals are shown.
Second, as noted in the main text, we found that sample splitting and refitting the model (see Appendix K.2) was somewhat unstable given different splits of the data. To illustrate this point, Figure A14 shows the 25th-75th percentile (and median) of the AMCEs estimated across twenty repetitions of splitting the data into halves and then using the refitting procedure described above. We address the problem of label switching using a permutation of labels that minimizes the average mean absolute error between all pairs of estimates; we find a permutation by randomly permuting the labels for a randomly chosen set of estimates and repeat this repeatedly until the average mean absolute error stabilizes.
While Figure A14 shows instability in some of the estimated AMCE, it broadly shows a similar result to that in the main text. For example, one group (Group 2 when $K=2$; Group 3 when $K=3$) shows a clear effect of country across most splits whereas one group (Group 1 when $K=2$ and Groups 1 and 2 when $K=2$) generally shows a large penalty for immigrants who entered without legal authorization.
Figure A14: The distribution of AMCE from a two-group and three- model with twenty random splits of the data. The interquartile range and median are shown.
Third, Figure 6 in the main text reports the average effect of changing some moderator from $x_{0}$ to $x_{1}$ on $\pi_{k}$, i.e.,
$\mathbb{E}\left[\pi_{k}(X_{ij}=x_{1},\bm{X}_{i,-j})-\pi_{k}(X_{ij}=x_{0},\bm{X}_{i,-j})\right].$ (A2)
Figure A15 considers the impact on the average absolute distance, i.e.
$\mathbb{E}\left[\left|\pi_{k}(X_{ij}=x_{1},\bm{X}_{i,-j})-\pi_{k}(X_{ij}=x_{0},\bm{X}_{i,-j})\right|\right],$ (A3)
to prevent positive and negative changes from canceling each other out. To interpret this quantity, Figure A15 also the absolute value of the difference reported in the main text, i.e., the absolute value of Equation A2 in a red *. Uncertainty is computed by drawing samples from the estimated asymptotic distribution of $\hat{\bm{\phi}}$, evaluating Equation A3 over those samples, and reporting the mean and $[0.025,0.975]$ percentile interval. Figure A15 shows that, for certain groups, some covariates show a small average effect but a larger average of absolute effects (e.g., with $K=3$, Group 2 and “Not Strong Republican” versus the baseline of “Strong Republican”) .
Figure A15: The average absolute effect of changing a moderator. The 2.5% to 97.5% percentile interval is shown.
Next, we discuss the two-factor interactions. The largest average marginal interaction effect (AMIE) was found between education and job in the three group analysis. This is visualized in Figure A16. The largest AMIE occurs between the levels of Teacher and High School and has magnitude of 0.0021.
Compared in magnitude to the AME, which for education was on average 0.111 and for job was on average 0.0237, this is clearly negligible. Given this, we have little hope of finding substantial higher-order interactions in this example.
If higher-order interactions were of interest, a pre-processing step to do some basic screening (see, e.g., Shi, Wang and Ding, 2023) might be implemented on the full dataset to a priori reduce the number of interactions considered. The sparsity inducing penalties of our method would then impose additional regularization.
Figure A16: The average marginal interaction effect between education and job.
Finally, we briefly remark upon choosing $K$ using an information criterion. While this works well in the simulated example (see Appendix K.3.1), we find less clear results on the full data. Table K.3.1 the results of optimizing the BIC over $\lambda$ for $K\in\{1,2,3,4\}$ as well as optimizing the AIC over $\lambda$. It shows that, if one uses the BIC, this suggests $K=1$. However, if one uses the AIC, this suggests $K=4$.
Table A3: Information criterion for different $K$